### v1

# Convergence Analysis of Natural Gradient Descent for Over-parameterized Physics-Informed Neural Networks

Abstract.
First-order methods, such as gradient descent (GD) and stochastic gradient descent (SGD) have been proven effective in training neural networks. In the setting of over-parameterization, there is a line of work demonstrating that randomly initialized (stochastic) gradient descent converges to a globally optimal solution at a linear convergence rate for the quadratic loss function. However, the learning rate of GD in training two-layer neural networks has a poor dependence on the sample size and the Gram matrix, resulting in a slow training process. In this paper, we show that for the $L^{2}$ regression problems, the learning rate can be improved from $\mathcal{O}(\lambda_{0}/n^{2})$ to $\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$, which implies that GD enjoys a faster convergence rate. Moreover, we further generalize the method for GD in training two-layer Physics-Informed Neural Networks (PINNs), showing a similar improvement for the learning rate. Although the improved learning rate depends mildly on the Gram matrix, we still need to set it small enough in practice due to the agnostic eigenvalues of the Gram matrix. More importantly, the convergence rate relies on the least eigenvalue of the Gram matrix, leading to slow convergence. In this work, we provide the convergence analysis of natural gradient descent (NGD) in training two-layer PINNs. We show that the learning rate can be $\mathcal{O}(1)$ and at this time, the convergence rate is independent of the Gram matrix.
* Corresponding author.
1. Introduction
In recent years, neural networks have achieved remarkable breakthroughs in the fields of image recognition [1], natural language procssing [2], reinforcement learning [3] and so on. Moreover, due to the flexibility and scalability of neural networks, researchers are paying much attention in exloring new methods involving neural networks for handling problems in scientific computing, where one long-standing and essential problem is solving partial differntial equations (PDEs) numerically. Classical numerical methods, such as finite difference, finite volume and finite elements methods, suffer the curse of dimensionality in solving high-dimensional PDEs. Due to this drawback, various methods involving neural networks have been proposed for solving different type PDEs [4, 5, 6, 7, 8]. Among them, the most representative approach is Physics-Informed Neural Networks (PINNs) [5]. In the framework of PINNs, one incorporate PDE constraints into the loss function and train the neural network with it. With the use of automatic differentiation, the neural network can be efficiently trained by first-order or second-order methods.
In the applications of neural networks, one inevitable issue is the selection of the optimization method. First-order methods, such as gradient descent (GD) and stochastic gradient descent (SGD), are widely used in optimizing neural networks as they only calculate the gradient and therefore are computationally efficient. Beside the first-order methods, there has been much interest in utilizing second-order optimization methods to speed up the training, not only for the regression problems [9] but also for the problems related to PDEs [4, 5].
For the convergence part of the optimization method, it has been shown that gradient descent algorithm can even achieve zero training loss under the setting of over-parametrization, which refers to a situation where a model has more parameters than necessary to fit the data [10, 11, 12, 13, 14, 15]. These works are based on the idea of neural tangent kernel (NTK), which shows that training multi-layer fully-connected neural networks via gradient descent is equivalent to performing a certain kernel method as the width of every layer goes to infinity. As for the finite width neural networks, with more refined analysis, it can be shown that the parameters are closed to the initialization throughout the entire training process when the width is large enough. It directly leads to the linear convergence for GD. Despite these beautiful convergence results, the learning rate depends on the sample size and the Gram matrix. Thus, it need to be sufficiently small to guarantee the convergence in practice. But doing so results in a slow training process. In contrast to the first-order methods, the second-order methods enjoy fast convergence due to the absence of requirements for the learning rate [16].
1.1. Contributions
The main contributions of our work can are summarized as follows:
  • For the $L^{2}$ regression problems, we demonstrate that the learning rate $\eta$ of gradient descent can be improved from $\mathcal{O}(\lambda_{0}/n^{2})$, as shown in [10], to $\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$, where $\bm{H}^{\infty}$ is the Gram matrix induced by the ReLU activation function and the random initialization, $\lambda_{0}$ is the least eigenvalue of $\bm{H}^{\infty}$. [17] has also shown the same improvement of the learning rate, but it requires that $m=\Omega(\frac{n^{6}}{\lambda_{0}^{4}\delta^{3}})$. Moreover in [18], the dependence on $n$, i.e. $\Omega(n^{6})$, is necessary and can not be improved due to the requirement of the proof method. Different from the method in [17], our method comes from a new recursion formula of gradient descent and it can be easily generalized to PINNs. Compared to [17], we only require that $m=\Omega(\frac{n^{4}}{\lambda_{0}^{4}}(\log(\frac{n}{\delta}))^{2})$.
  • For the PINNs, we simultaneously improve both the learning rate $\eta$ of gradient descent and the requirement of the width $m$. The improvements rely on a new recursion formula of gradient descent, which is similar to that for regression problems. Specifically, we can improve the learning rate $\eta=\mathcal{O}(\lambda_{0})$ required in [19] to $\eta=\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$ and the requirement for the width $m$, i.e. $m=\tilde{\Omega}(\frac{(n_{1}+n_{2})^{2}}{\lambda_{0}^{4}\delta^{3}})$, can be improved to $m=\tilde{\Omega}(\frac{1}{\lambda_{0}^{4}}(\log(\frac{n_{1}+n_{2}}{\delta})))$.
  • We provide the convergence results for natural gradient descent (NGD) in training over-parameterized two-layer PINNs with $\text{ReLU}^{3}$ activation functions and smooth activation functions. Compared to gradient descent, the convergence rate of NGD is independent of $\lambda_{0}$, which results in a faster convergence. Meanwhile, we show that the learning rate can be even $\mathcal{O}(1)$. When the activation function is smooth, NGD enjoys quadratic convergence rate.
1.2. Related Works
First-order methods. There are mainly two ways to study the optimization of neural networks and try to understand why first-order methods can find a global minimum. One way is to analyse the optimization landscape [15, 20]. It has been shown that gradient descent can find a global minimum in polynomial time if the optimization landscape has some good geometric properties. However, some unrealistic assumptions in these works make it difficult to generalize to practical neural networks. Another way to understand the optimization of neural networks is to analyze the optimization dynamics of first-order methods. For the two-layer ReLU neural networks, as shown in [10], as long as the width $m$ is large enough and no two inputs are parallel, randomly initialized gradient descent converges to a globally optimal solution at a linear rate. Later, the results have been generalized to deep neural networks with smooth activation functions [11]. Both rely on that the Gram matrices are stable throughout the training process, which leads to the convergence to the global minimum. Besides regression and classification problems, [19] demonstrated the convergence of the gradient descent for two-layer PINNs by the similar analysis of optimization dynamics. However, both [10] and [19] require that the learning rate be sufficiently small to converge. In this work, we conduct a refined analysis of gradient descent for $L^{2}$ regression problems and PINNs, leading to a milder requirement for the learning rate.
Second-order methods. Although second-order methods possess better convergence rate, they are rarely used in training deep neural networks due to the prohibitive computational cost. As a variant of the original Gauss-Newton method, natural gradient descent (NGD) is more efficient in practice. Meanwhile, as shown in [21] and [16], it also enjoys faster convergence rate for the $L^{2}$ regression problems than of gradient descent. In this paper, we provide the convergence analysis for NGD in training two-layer PINNs.
1.3. Notations
We denote $[n]=\{1,2,\cdots,n\}$ for $n\in\mathbb{N}$. Given a set $S$, we denote the uniform distribution on $S$ by $Unif\{S\}$. We use $I\{E\}$ to denote the indicator function of the event $E$. For variables $a$ and $b$, $a\lesssim b$ is equivalent to $a\leq Cb$ for some universal constant $C$. For two functions $f_{1}(n)$ and $f_{2}(n)$, the relation $f_{1}(n)=\mathcal{O}(f_{2}(n))$ is equivalent to $f_{1}(n)\leq Cf_{2}(n)$ for some universal constant $C$. And similarly, $f_{1}(n)=\Omega(f_{2}(n))$ is equivalent to $f_{1}(n)\geq Cf_{2}(n)$. Throughout the paper, we use boldface denote vectors. Given $x_{1},\cdots,x_{d}\in\mathbb{R}$, we use $(x_{1},\cdots,x_{d})$ or $[x_{1},\cdots,x_{d}]$ to denote a row vector with $i$-th component $x_{i}$ for $i\in[d]$ and then $(x_{1},\cdots,x_{d})^{T}\in\mathbb{R}^{d}$ is a column vector.
1.4. Organization of this Paper
We present the improvements of the learning rate of gradient descent for $L^{2}$ regression problems and PINNs in Section 2 and Section 3 respectively. In Section 4, we show the convergence of natural gradient descent for training PINNs with $\text{ReLU}^{3}$ activation functions and smooth activation functions. We conclude in Section 5 and the detailed proofs are put in the Appendix for readability and brevity.
2. Improved Learning Rate of Gradient Descent for $L^{2}$ Regression Problems
2.1. Problem Setup
In this section, we consider a two-layer neural network $f$ with the following form.
$f(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\tilde{\bm{x}}),$ (1)
where $\bm{x}\in\mathbb{R}^{d}$ is the input, $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+1}$, $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T},\bm{a}=(a_{1},\cdots,a_{m})^{T}$, for $r\in[m]$, $\bm{w}_{r}\in\mathbb{R}^{d+1}$ is the weight vector of the first layer, $a_{r}\in\mathbb{R}$ is the output weight and $\sigma(\cdot)$ is the ReLU activation function. Different from the setup in [10], the two-layer neural network that we consider has bias term. In the following, we assume that each input vector $\bm{x}\in\mathbb{R}^{d}$ has been augmented to $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+1}$. With a little abuse of notation, we write $\bm{x}$ for $\tilde{\bm{x}}$ and write $f$ as
$f(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\bm{x}).$
For the $L^{2}$ regression problem, given training data set $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, the aim is to minimize the loss function
$L(\bm{w},\bm{a}):=\sum\limits_{i=1}^{n}\frac{1}{2}(f(\bm{x}_{i};\bm{w},\bm{a})-y_{i})^{2}.$ (2)
Before training, we initialize the first layer vector $\bm{w}_{r}(0)\sim\mathcal{N}(\bm{0},\bm{I})$ and output weight $a_{r}\sim Unif(\{-1,1\})$ for $r\in[m]$. In the training process, we fix the second layer and optimize the first layer by gradient descent (GD), i.e., we update the weights by the following formulation.
$\bm{w}(k+1)=\bm{w}(k)-\eta\frac{\partial L(\bm{w}(k))}{\partial\bm{w}},$ (3)
where $\eta>0$ is the learning rate, $k\in\mathbb{N}$ and $L(\bm{w}(k))$ is an abbreviation of $L(\bm{w}(k),\bm{a})$. From the form of $L(\bm{w})$, we can deduce that
$\frac{\partial L(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{m}}\sum\limits_{i=1}^{n}(f(\bm{x}_{i};\bm{w},\bm{a})-y_{i})\bm{x}_{i}I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}.$ (4)
We denote $u_{i}(k)=f(\bm{x}_{i};\bm{w}(k),\bm{a})$ the prediction on input $\bm{x}_{i}$ at $k$-th iteration and we let $\bm{u}(k)=(u_{1}(k),\cdots,u_{n}(k))^{T}\in\mathbb{R}^{n}$ be the prediction vector at $k$-th iteration. For the labels, we define $\bm{y}:=(y_{1},\cdots,y_{n})^{T}\in\mathbb{R}^{n}$.
According to [10], in the continuous setting, the dynamics of predictions can be written as
$\frac{d}{dt}\bm{u}(t)=\bm{H}(t)(\bm{y}-\bm{u}(t)),$ (5)
where $H(t)\in\mathbb{R}^{n\times n}$ is the Gram matrix at time $t$ with $(i,j)$-entry
$H_{ij}(t)=\frac{1}{m}\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}_{r}(t)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(t)^{T}\bm{x}_{j}\geq 0\right\}.$ (6)
In the setting of over-parameterization and randomly initialization, [10] has shown that (1) $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}=\mathcal{O}(\sqrt{1/m})$, where $\bm{H}^{\infty}$ is the Gram matrix induced by the initialization with $(i,j)$-th entry
$H_{ij}^{\infty}=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}^{T}\bm{x}_{i}\geq 0,\bm{w}^{T}\bm{x}_{j}\geq 0\right\}\right]$ (7)
and (2) $\|\bm{H}(t)-\bm{H}(0)\|_{2}=\mathcal{O}(\sqrt{1/m})$ for all $t>0$. Therefore, as $m\to\infty$, the dynamics of the predictions are charactered by $H^{\infty}$, leading to the linear convergence.
2.2. Main Results
To simplify the analysis, we make the following assumptions on the training data.
Assumption 1.
For $i\in[n]$, $\|\bm{x}_{i}\|_{2}\leq\sqrt{2}$ and $|y_{i}|\leq 1$, where $\bm{x}_{i}\in\mathbb{R}^{d+1}$ is the augmented input.
Assumption 2.
No two samples in $\{\bm{x}_{i}\}_{i=1}^{n}$ are parallel, i.e., for any $\bm{x}_{t},\bm{x}_{s}\in\{\bm{x}_{i}\}_{i=1}^{n}$ and any $\alpha\in\mathbb{R}$, we have $\bm{x}_{t}\neq\bm{x}_{s}$.
Since inputs are all augmented, Assumption 2 is equivalent to that no two samples in $\{\bm{x}_{i}\}_{i=1}^{n}$ are equal, which holds naturally. Under Assumption 2, Theorem 3.1 in [10] implies that $\lambda_{0}:=\lambda_{min}(\bm{H}^{\infty})>0$, which is crucial in the convergence analysis.
As stated before, there are two improtant facts that determine the optimization dynamics, one is that at initialization $\bm{H}(0)$ is closed to $\bm{H}^{\infty}$ and another is that $\bm{H}(k)$ does not go far away from the initialization $\bm{H}(0)$ for all $k\in\mathbb{N}$. These facts are supported by the following two lemmas.
Lemma 1 (Lemma 3.1 in [10]).
If $m=\Omega\left(\frac{n^{2}}{\lambda_{0}^{2}}\log\left(\frac{n}{\delta}\right)\right)$, we have with probability at least $1-\delta$, $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}\leq\frac{\lambda_{0}}{4}$ and $\lambda_{min}(\bm{H}(0))\geq\frac{3\lambda_{0}}{4}$.
Lemma 2.
Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d. generated from $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-n^{2}e^{-mR/4}$, the following holds. For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+1}$ that satisfy for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then the matrix $\bm{H}(\bm{w})\in\mathbb{R}^{n\times n}$ defined by
$H(\bm{w})_{ij}=\frac{1}{m}\bm{x}_{i}^{T}\bm{x}_{j}\sum\limits_{r=1}^{m}I\left\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\right\}.$
satisfies
$\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}<4nR.$ (8)
Remark 1.
Although Lemma 2 looks almost same to Lemma 3.1 in [18], the proof of Lemma 3.1 in [18] lacks an important part, i.e., the proof of the measurability for the related variable. In fact, Lemma 2 ensures us to bound the change of $\bm{H}(\bm{w})$ when $\bm{w}$ in a small ball. However, the set consisting of vectors in the small ball is uncountable, thus the related variable may not be measurable due to the discontinuity of the indicator function appeared in the definition of $\bm{H}(\bm{w})$. In the proof, we borrow the concept of pointwise measurability in the empirical process (see Chapter 8 in [22]) to bridge the gap. Then we can bound the supremum with a new random variable. By applying a concentration inequality to this new random variable, we can reach the conclusion.
Abount the convergence of gradient descent, [10] has demonstrated that if the learning rate $\eta=\mathcal{O}(\lambda_{0}/n^{2})$, then randomly initialized gradient descent converges to a globally optimal solution at a linear convergence rate when $m$ is large enough. The requirement of $\eta$ is from the decomposition for the residual in the $(k+1)$-th iteration, i.e.,
$\bm{y}-\bm{u}(k+1)=\bm{y}-\bm{u}(k)-(\bm{u}(k+1)-\bm{u}(k)).$ (9)
Instead of decomposing the residual into the two terms as above, we write it as follows, which serves as a recursion formula.
Lemma 3.
For all $k\in\mathbb{N}$, we have
$\bm{y}-\bm{u}(k+1)=(\bm{I}-\eta H(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k),$ (10)
where $\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n}(k))^{T}\in\mathbb{R}^{n}$ with $i$-th
$I_{1}^{i}(k):=u_{i}(k+1)-u_{i}(k)-\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle.$ (11)
After [10], [17] improved the network size $m$ to $m=\mathcal{O}(\lambda_{0}^{-4}n^{4}\log^{3}(n/\delta)\log(m/\delta))$, but it still requires $\eta=\mathcal{O}(\lambda_{0}/n^{2})$. Another work [17] provided a refined analysis over [10], which shows that $\eta=\frac{1}{C_{1}\|\bm{H}^{\infty}\|_{2}}$ is enough for the convergence rate $1-\frac{\lambda_{0}C_{2}}{C_{1}\|\bm{H}{\infty}\|_{2}}$, but the constants $C_{1},C_{2}$ may depend on the parameters $\lambda_{0},n$ and $\delta$. It requires the network size $m$ to be $\Omega(\lambda_{0}^{-4}n^{6}\delta^{-3})$ to make $C_{1},C_{2}$ independent of $\lambda_{0},n$ and $\delta$. In this paper, we improve both the learning rate and network size and the main reslut is the following theorem.
Theorem 1.
Under Assumption 1 and Assumption 2, if we set the number of hidden nodes $m=\Omega\left(\frac{n^{4}}{\lambda_{0}^{4}}\log\left(\frac{n}{\delta}\right)\right)$ and the learning rate $\eta=\mathcal{O}\left(\frac{1}{\|\bm{H}^{\infty}\|_{2}}\right)$, then with probability at least $1-\delta$ over the random initialization, the gradient descent algorithm satisfies
$\|\bm{y}-\bm{u}(k)\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{k}\|\bm{y}-\bm{u}(0)\|_{2}^{2}$ (12)
for all $k\in\mathbb{N}$.
Remark 2.
Our proof method requires that $\bm{I}-\eta\bm{H}(k)$ is positive definite for all $k\in\mathbb{N}$. As $\bm{H}(k)$ is closed to $\bm{H}^{\infty}$, the requirement is satisfied if $\eta\|\bm{H}^{\infty}\|_{2}\lesssim 1$. Note that $\|\bm{H}^{\infty}\|_{2}\leq n$, it is sufficient to set $\eta=\mathcal{O}(1/n)$, which results in an improvement of $\mathcal{O}(\lambda_{0}/n)$.
3. Improved Learning Rate of Gradient Descent for Two-Layer Physics-Informed Neural Networks
3.1. Problem Setup
In this section, we consider the same setup as [19], i.e., we focus on the PDE with the following form.
$&\frac{\partial u}{\partial x_{0}}(\bm{x})-\sum\limits_{i=1}^{d}\frac{\partial^{2}u}{\partial x_{i}^{2}}(\bm{x})=f(\bm{x}),\,\bm{x}\in(0,T)\times\Omega,\\ &u(\bm{x})=g(\bm{x}),\ \bm{x}\in\{0\}\times\Omega\cup[0,T]\times\partial\Omega,\end{aligned}\right.$ (13)
where $\bm{x}=(x_{0},x_{1},\cdots,x_{d})^{T}\in\mathbb{R}^{d+1}$ and $x_{0}\in[0,T]$ is the time variable. In the following, we assume that $\|\bm{x}\|_{2}\leq 1$ for $\bm{x}\in[0,T]\times\bar{\Omega}$ and $f,g$ are bounded continuous functions.
Moreover, we consider a two-layer neural network of the following form.
$\phi(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\tilde{\bm{x}}),$ (14)
where $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T}\in\mathbb{R}^{m(d+2)}$, $\bm{a}=(a_{1},\cdots,a_{m})^{T}\in\mathbb{R}^{m}$ and for $r\in[m]$, $\bm{w}_{r}\in\mathbb{R}^{d+2}$ is the weight vector of the first layer, $a_{r}$ is the output weight and $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function. Similar as that in Section 2, $\tilde{\bm{x}}=(\bm{x}^{T},1)^{T}\in\mathbb{R}^{d+2}$ is the augmented vector from $\bm{x}$ and in the following, we write $\bm{x}$ for $\tilde{\bm{x}}$ for brevity.
In the framework of PINNs, given training samples $\{\bm{x}_{p}\}_{p=1}^{n_{1}}$ and $\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ that are from interior and boundary respectively, we aim to minimize the following empirical loss function.
$L(\bm{w},\bm{a})$ $:=\sum\limits_{p=1}^{n_{1}}\frac{1}{2n_{1}}\left(\frac{\partial\phi}{\partial x_{0}}(\bm{x}_{p};\bm{w},\bm{a})-\sum\limits_{i=1}^{d}\frac{\partial^{2}\phi}{\partial x_{i}^{2}}(\bm{x}_{p};\bm{w},\bm{a})-f(\bm{x}_{p})\right)^{2}$ (15)
$\quad+\sum\limits_{j=1}^{n_{2}}\frac{1}{2n_{2}}\left(\phi(\bm{y}_{j};\bm{w},\bm{a})-g(\bm{y}_{j})\right)^{2}.$
Similar as that for the $L^{2}$ regression problems, we initialize the first layer vector $\bm{w}_{r}(0)\sim\mathcal{N}(\bm{0},\bm{I})$, output weight $a_{r}\sim Unif(\{-1,1\})$ for $r\in[m]$ and fix the output weights. Then the gradient descent updates the hidden weights by the following formulations:
$\bm{w}_{r}(k+1)$ $=\bm{w}_{r}(k)-\eta\frac{\partial L(\bm{w}(k),\bm{a})}{\partial\bm{w}_{r}}$ (16)
for all $r\in[m]$ and $k\in\mathbb{N}$, where $\eta>0$ is the learning rate. For brevity, we write $L(\bm{w})$ for $L(\bm{w},\bm{a})$.
To simplify the notations, for the residuals of interior and boundary, we denote them by $s_{p}(\bm{w})$ and $h_{j}(\bm{w})$ respectively, i.e.,
$s_{p}(\bm{w})=\frac{1}{\sqrt{n_{1}}}\left(\frac{\partial u}{\partial x_{0}}(\bm{x}_{p};\bm{w})-\sum\limits_{i=1}^{d}\frac{\partial^{2}u}{\partial x_{i}^{2}}(\bm{x}_{p};\bm{w})-f(\bm{x}_{p})\right)$ (17)
and
$h_{j}(\bm{w})=\frac{1}{\sqrt{n_{2}}}(u(\bm{y}_{j};\bm{w})-g(\bm{y}_{j})).$ (18)
Then the empirical loss function can be written as
$L(\bm{w})=\frac{1}{2}\left(\|\bm{s}(\bm{w})\|_{2}^{2}+\|\bm{h}(\bm{w})\|_{2}^{2}\right),$ (19)
where
$\bm{s}(\bm{w})=(s_{1}(\bm{w}),\cdots,s_{n_{1}}(\bm{w}))^{T}\in\mathbb{R}^{n_{1}}$ (20)
and
$\bm{h}(\bm{w})=(h_{1}(\bm{w}),\cdots,h_{n_{2}}(\bm{w}))^{T}\in\mathbb{R}^{n_{2}}.$ (21)
At this time, we have
$\frac{\partial L(\bm{w})}{\partial\bm{w}_{r}}=\sum\limits_{p=1}^{n_{1}}s_{p}(\bm{w})\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}+\sum\limits_{j=1}^{n_{2}}h_{j}(\bm{w})\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}$ (22)
and the Gram matrix $\bm{H}(\bm{w})$ is defined as
$\bm{H}(\bm{w})=\bm{D}^{T}\bm{D},\ \bm{D}:=\left[\frac{\partial s_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial s_{n_{1}}(\bm{w})}{\partial\bm{w}},\frac{\partial h_{1}(\bm{w})}{\partial\bm{w}},\cdots,\frac{\partial h_{n_{2}}(\bm{w})}{\partial\bm{w}}\right].$ (23)
3.2. Main Results
First, we make the following assumptions about the training samples, which are similar to those for the $L^{2}$ regression problems.
Assumption 3.
For $p\in[n_{1}]$ and $j\in[n_{2}]$, $\|\bm{x}_{p}\|_{2}\leq\sqrt{2},\|\bm{y}_{j}\|_{2}\leq\sqrt{2}$, where all inputs have been augmented.
Assumption 4.
No two samples in $\{\bm{x}_{p}\}_{p=1}^{n_{1}}\cup\{\bm{y}_{j}\}_{j=1}^{n_{2}}$ are parallel.
Under Assumption 4, Lemma 3.3 in [19] implies that the Gram matrix $\bm{H}^{\infty}:=\mathbb{E}_{\bm{w}\sim\mathcal{N}(\bm{0},\bm{I})}\bm{H}(\bm{w})$ is strictly positive definite and we let $\lambda_{0}=\lambda_{min}(\bm{H}^{\infty})$. Similarly, $\bm{H}^{\infty}$ plays an important role in the optimization process. The following two lemmas indicate that $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}=\mathcal{O}(1/\sqrt{m})$ and $\|\bm{H}(k)-\bm{H}(0)\|_{2}=\mathcal{O}(1/\sqrt{m})$, leading to the linear convergence.
Lemma 4.
If $m=\Omega\left(\frac{d^{4}}{\lambda_{0}^{2}}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$, we have that with probability at least $1-\delta$, $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}\leq\frac{\lambda_{0}}{4}$ and $\lambda_{min}(\bm{H}(0))\geq\frac{3}{4}\lambda_{0}$.
Lemma 5.
Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d. generated from $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-\delta-(n_{1}+n_{2})^{2}e^{-mR}$, the following holds. For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+1}$ that satisfy for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then
$\|H(\bm{w})-H(0)\|_{F}<CM^{2}R,$ (24)
where $M=2(d+2)\log(2m(d+2)/\delta)$ and $C$ is a universal constant.
In [19], the decomposition for the residual in the $(k+1)$-th iteration is same as the one in [10], i.e.,
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\left(\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}-\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right),$ (25)
which leads to the requirement $\eta=\mathcal{O}(\lambda_{0})$. Instead, we can generalize easily the method used in the $L^{2}$ regression problems to PINNs and obtain the following recursion formula.
Lemma 6.
For all $k\in\mathbb{N}$, we have
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k),$ (26)
where $\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n_{1}}(k),I_{1}^{n_{1}+1}(k),\cdots,I_{1}^{n_{1}+n_{2}}(k))^{T}\in\mathbb{R}^{n_{1}+n_{2}}$ and for $p\in[n_{1}]$,
$I_{1}^{p}(k)=s_{p}(k+1)-s_{p}(k)-\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle,$ (27)
for $j\in[n_{2}]$,
$I_{1}^{n_{1}+j}(k)=h_{j}(k+1)-h_{j}(k)-\left\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle.$ (28)
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
By using the recursion formula, we arrive at our main result.
Theorem 2.
Under Assumption 3 and Assumption 4, if we set the number of hidden nodes
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
and the learning rate $\eta=\mathcal{O}\left(\frac{1}{\|\bm{H}^{\infty}\|_{2}}\right)$, then with probability at least $1-\delta$ over the random initialization, the gradient descent algorithm satisfies
$\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{k}\left\|\begin{bmatrix}\bm{s}(0)\\ \bm{h}(0)\end{bmatrix}\right\|_{2}^{2}$ (29)
for all $k\in\mathbb{N}$.
Remark 3.
It may be confused that [19] has uesd the same method in [10], but it only requires $\eta=\mathcal{O}(\lambda_{0})$. Actually, it is due to that the loss function of PINN has been normalized. If we let $n_{1}=n_{2}=n$ and $\widetilde{\bm{H}}^{\infty}$ be the Gram matrix induced by unnormalized loss function of PINN, then $\lambda_{min}(\bm{H}^{\infty})=\lambda_{min}(\widetilde{\bm{H}}^{\infty})/n$, leading to the convergence rate similar to that of regression problem. Due to the normalization of loss function, $\|\bm{H}^{\infty}\|_{2}$ is independent of the sample size $n$, but may depends on the dimension $d$.
4. Convergence of Natural Gradient Descent for Two-Layer Physics-Informed Neural Networks
4.1. Problem Setup
Although we have improved the learning rates of gradient descent for $L^{2}$ regression problems and PINNs, one needs to set the learning rates to be small enough due to the agnostic $\|\bm{H}^{\infty}\|_{2}$. For instance, we may need to set it to be $\mathcal{O}(1/n)$ for the $L^{2}$ regression problem. Moreover, the convergence rate $1-\frac{\eta\lambda_{0}}{2}$ also depends on $\lambda_{0}$, which may be slow with small $\lambda_{0}$. [21] and [16] have provided the convergence results for natural gradient descent (NGD) in training over-parameterized two-layer neural networks for $L^{2}$ regression problems. They showed that the maximal learning rate can be $\mathcal{O}(1)$ and the convergence rate is independent of $\lambda_{0}$, which result in a faster convergence rate. In the section, we conduct the convergence analysis of NGD for PINNs and show that NGD indeed leads to an even faster convergence rate.
We consider the same setup as Section 3 and aim to minimize the following empirical loss function via NGD.
$L(\bm{w}):=\frac{1}{2}\left(\|\bm{s}(\bm{w})\|_{2}^{2}+\|\bm{h}(\bm{w})\|_{2}^{2}\right).$ (30)
The NGD gives the following update rule:
$\bm{w}(k+1)=\bm{w}(k)-\eta\bm{J}(k)^{T}\left(\bm{J}(k)\bm{J}(k)^{T}\right)^{-1}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},$ (31)
where $\bm{J}(k)=\left(\bm{J}_{1}(k)^{T},\cdots,\bm{J}_{n_{1}+n_{2}}(k)^{T}\right)^{T}\in\mathbb{R}^{(n_{1}+n_{2})\times m(d+2)}$ is the Jacobian matrix for the whole dataset and $\eta>0$ is the learning rate. Specifically, for $p\in[n_{1}]$,
$\bm{J}_{p}(k)=\left[\left(\frac{\partial s_{p}(k)}{\partial\bm{w}_{1}}\right)^{T},\cdots,\left(\frac{\partial s_{p}(k)}{\partial\bm{w}_{m}}\right)^{T}\right]\in\mathbb{R}^{1\times m(d+2)}$ (32)
and for $j\in[n_{2}]$,
$\bm{J}_{n_{1}+j}(k)=\left[\left(\frac{\partial h_{j}(k)}{\partial\bm{w}_{1}}\right)^{T},\cdots,\left(\frac{\partial h_{j}(k)}{\partial\bm{w}_{m}}\right)^{T}\right]\in\mathbb{R}^{1\times m(d+2)}.$ (33)
For the activation function of the two-layer neural network
$\phi(\bm{x};\bm{w},\bm{a})=\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}^{T}\bm{x}),$ (34)
we consider both settings that $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function or a smooth activation function that satisfies the following assumption. From Lemma 3.3 in [19] and Lemma 2 in [23], we know that $\bm{H}^{\infty}$ is strictly positive definite in both settings and we let $\lambda_{0}=\lambda_{min}(\bm{H}^{\infty})$.
Assumption 5.
There exists a constant $c>0$ such that $|\sigma(0)|\leq c$ and for any $z,z^{{}^{\prime}}\in\mathbb{R}$,
$|\sigma^{(k)}(z)-\sigma^{(k)}(z^{{}^{\prime}})|\leq c|z-z^{{}^{\prime}}|,$ (35)
where $k\in\{0,1,2,3\}$. Moreover, $\sigma(\cdot)$ is analytic and is not a polynomial function.
Unlike the approach for gradient descent, [21] and [16] focus on the change of the Jacobian matrix for NGD, rather than focusing on the Gram matrix. More precisely, they show that $\bm{J}(\bm{w})$ is stable with respect to $\bm{w}$, where $\bm{J}(\bm{w})$ is the Jacobian matrix with weight vector $\bm{w}=(\bm{w}_{1}^{T},\cdots,\bm{w}_{m}^{T})^{T}$. However, this approach does not work for PINNs. Instead, we consider the stability of $\bm{J}(\bm{w})$ with respect to every single weight vector $\bm{w}_{r}$.
Lemma 7 (Stable Jacobian).
Let $R\in(0,1]$, if $\bm{w}_{1}(0),\cdots,\bm{w}_{m}(0)$ are i.i.d. generated $\mathcal{N}(\bm{0},\bm{I})$, then with probability at least $1-P(\delta,m,R)$ the following holds. For any set of weight vectors $\bm{w}_{1},\cdots,\bm{w}_{m}\in\mathbb{R}^{d+2}$ that satsify for any $r\in[m]$, $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, then
(1) when $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function, we have that
$\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}\leq CM\sqrt{R},$ (36)
where $C$ is a universal constant, $M=2(d+2)\log(2m(d+2)/\delta)$ and
$P(\delta,m,R)=1-\delta-n_{1}e^{-mR};$ (37)
(2) when $\sigma(\cdot)$ satisfies Assumption 5, we have that
$\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}\leq CdR$ (38)
for $m\geq\log^{2}(1/\delta)$, where $C$ is a universal constant and $P(\delta,m,R)=\delta$.
With the stability of Jacobian matrix, we can derive the following main result.
Theorem 3 (Convergence of NGD).
Let $L(k)=L(\bm{w}(k))$, then the following conclusions hold.
(1) When $\sigma(\cdot)$ is the $\text{ReLU}^{3}$ activation function, under Assumption 4, we set
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
and $\eta\in(0,1)$, then with probability at least $1-\delta$ over the random initialization for all $k\in\mathbb{N}$
$L(k)\leq(1-\eta)^{k}L(0).$ (39)
(2) When $\sigma(\cdot)$ satisfies Assumption 5, under Assumption 4, we set
$m=\Omega\left(\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
and $\eta\in(0,1)$, then with probability at least $1-\delta$ over the random initialization for all $k\in\mathbb{N}$
$L(k)\leq(1-\eta)^{k}L(0).$ (40)
In fact, NGD can enjoy a second-order convergence rate when $\sigma(\cdot)$ satisfies Assumption 5.
Corollary 1.
Under the setting of Theorem 3(2), in addition to (40), we also have that
$\left\|\begin{pmatrix}\bm{s}(t+1)\\ \bm{h}(t+1)\end{pmatrix}\right\|_{2}\leq\frac{CB^{4}}{\sqrt{m\lambda_{0}^{3}}}\left\|\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}\right\|_{2}^{2},$
where $C$ is a universal constant and $B=\sqrt{2(d+2)\log(2m(d+2)/\delta)}+1$.
5. Conclusion and Discussion
In this paper, we have improved the learning rate of gradient descent for both $L^{2}$ regression problems and PINNs, indicating that gradient descent enjoys better convergence. Furthermore, we demonstrate that natural gradient descent can find the global optima of two-layer PINNs with $\text{ReLU}^{3}$ or smooth activation functions for a class of second-order linear PDEs. Compared to gradient descent, natural gradient descent possesses faster convergence rate and the maximal learning rate is $\mathcal{O}(1)$. Despite this, natural gradient descent is quite expensive in terms of cumputation and memory in training neural networks. Therefore, several cost-effective variants have been proposed, such as K-FAC [9] and mini-batch natural gradient descent [16]. It would be interesting to investigate the convergence of these methods for PINNs. Additionally, generalizing the convergence analysis to deep neural networks is an important direction for future work.
Appendix
6. Proof of Section 2
Before the proofs, we first define the event
$A_{ir}:=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}(0)\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{i}\geq 0\}\neq I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0\}\}$ (41)
for all $i\in[n]$.
Note that the event happens if and only if $|\bm{w}_{r}(0)^{T}\bm{x}_{i}|<R$, thus by the anti-concentration inequality of Gaussian distribution, we have
$P(A_{ir})=P_{z\sim\mathcal{N}(0,\|\bm{x}_{i}\|_{2}^{2})}(|z|<R)=P_{z\sim\mathcal{N}(0,1)}\left(|z|<\frac{R}{\|\bm{x}_{i}\|_{2}}\right)\leq\frac{2R}{\sqrt{2\pi}\|\bm{x}_{i}\|_{2}}\leq\frac{2R}{\sqrt{2\pi}},$ (42)
as $\|\bm{x}_{i}\|_{2}\geq 1$ for all $i\in[n]$.
6.1. Proof of Lemma 2
Proof.
Recall that
$\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(\bm{w})_{ij}-H(0)_{ij}|^{2}$ (43)
$=\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}(\bm{x}_{i}^{T}\bm{x}_{j})^{2}\left(\frac{1}{m}\sum\limits_{r=1}^{m}(I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\})\right)^{2}.$
To prove the measurability, we will show that $I\left\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\right\}$ can be approximated by the sequence $\{I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0,\tilde{\bm{w}}_{k}^{T}\bm{x}_{j}\geq 0\}\}_{k\in\mathbb{N}}$, where $\tilde{\bm{w}}_{k}$ is a rational vector in $\mathbb{Q}^{d+1}$ for each $k\in\mathbb{N}$, with $\mathbb{Q}$ denoting the set of all rational numbers. Therefore, taking the supremum of $\bm{w}$ within the ball is equivalent to taking the supremum over all rational vectors within this ball, which implies that the random variable is measurable.
We first focus on $I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}$ and let $\bm{w}_{r}=(\bm{w}_{r1}^{T},w_{r0})^{T},\bm{x}_{i}=(\bm{x}_{i1}^{T},1)^{T}$ with $w_{r0}\in\mathbb{R}$ and $\bm{w}_{r1},\bm{x}_{i1}\in\mathbb{R}^{d}$. For each $k\in\mathbb{N}$, we can take $\bm{\beta}_{k}\in\mathbb{Q}^{d}$ such that $\|\bm{\beta}_{k}-\bm{w}_{r1}\|_{2}\leq\frac{1}{2k}$. Then choose $t_{k}\in\mathbb{Q}$ such that $t\in(w_{r0}+\frac{1}{2k},w_{r0}+\frac{1}{k}]$ and let $\tilde{\bm{w}}_{k}=(\bm{\beta}_{k}^{T},t_{k})^{T}$. From the construction, we have
$I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0\}=I\{\bm{\beta}_{k}^{T}\bm{x}_{i1}+t_{k}\geq 0\}=I\{\bm{w}_{r}^{T}\bm{x}_{i}+t_{k}-w_{r0}+(\bm{\beta}_{k}-\bm{w}_{r1})^{T}\bm{x}_{i1}\geq 0\},$
as $\bm{w}_{r}^{T}\bm{x}_{i}=\bm{w}_{r1}^{T}\bm{x}_{i1}+w_{r0}$. Define $r_{k}=t_{k}-w_{r0}+(\beta_{k}-\bm{w}_{r1})^{T}\bm{x}_{i1}$, then $r_{k}>0$ and $r_{k}\to 0$, as $|(\beta_{k}-\bm{w}_{r1})^{T}\bm{x}_{i1}|\leq\|\beta_{k}-\bm{w}_{r1}\|_{2}\leq\frac{1}{2k}$ and $t-w_{r0}\in(\frac{1}{2k},\frac{1}{k}]$. Since the function $u\rightarrow I\{u\geq x\}$ is right-continuous for any $x\in\mathbb{R}$. Thus $I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0\}=I\{\bm{w}_{k}^{T}\bm{x}_{i}+r_{k}\geq 0\}\to I\{\bm{w}_{k}^{T}\bm{x}_{i}\geq 0\}$ as $k\to\infty$. Meanwhile, when $k$ is large enough, $\tilde{\bm{w}}_{k}$ is in the small ball centered at $\bm{w}_{r}(0)$, i.e., $\|\tilde{\bm{w}}_{k}-\bm{w}_{r}(0)\|_{2}<R$. Similarly, we have that $I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{j}\geq 0\}\to I\{\bm{w}_{k}^{T}\bm{x}_{j}\geq 0\}$ as $k\to\infty$. Thus, we can deduce that $I\{\tilde{\bm{w}}_{k}^{T}\bm{x}_{i}\geq 0,\tilde{\bm{w}}_{k}^{T}\bm{x}_{j}\geq 0\}\to I\{\bm{w}_{k}^{T}\bm{x}_{i}\geq 0,\bm{w}_{k}^{T}\bm{x}_{j}\geq 0\}$ as $k\to\infty$.
Therefore,
$\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\\ \cdots\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(\bm{w})_{ij}-H(0)_{ij}|^{2}=\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\bm{w}_{1}\in\mathbb{Q}^{d+1}\\ \cdots\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R,\bm{w}_{m}\in\mathbb{Q}^{d+1}\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(\bm{w})_{ij}-H(0)_{ij}|^{2},$
which leads to the measurability.
From the definition of $A_{ir}$, i.e., (41), and the fact $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, we can deduce that
$|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0\}|\leq I\{A_{ir}\}.$
Thus, when both $A_{ir}$ and $A_{jr}$ do not happen, we have
$|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\}|$
$\leq|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0\geq 0\}|+|I\{\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\geq 0\}|$
$\leq I\{A_{ir}\}+I\{A_{jr}\}=0,$
which imples
$|I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\}|\leq I\{A_{ir}\lor A_{jr}\}.$ (44)
Therefore, combining (43) and (44) yields that
$\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\\ \cdots,\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}|H(w)_{ij}-H(0)_{ij}|^{2}$ (45)
$\leq 4\sup\limits_{\begin{subarray}{c}\\ \|\bm{w}_{1}-\bm{w}_{1}(0)\|_{2}<R,\\ \cdots,\\ \|\bm{w}_{m}-\bm{w}_{m}(0)\|_{2}<R\end{subarray}}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\left(\frac{1}{m}\sum\limits_{r=1}^{m}(I\{\bm{w}_{r}^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{i}\geq 0,\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\})\right)^{2}$
$\leq 4\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}\left(\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}\right)^{2}.$
It remains only to bound the term $\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}$. Note that
$\mathbb{E}[I\{A_{ir}\lor A_{jr}\}]\leq P(A_{ir})+P(A_{jr})\leq\frac{4R}{\sqrt{2\pi}}$
and then
$Var(I\{A_{ir}\lor A_{jr}\})\leq\mathbb{E}[I\{A_{ir}\lor A_{jr}\}\leq\frac{4R}{\sqrt{2\pi}}.$
Thus, applying Bernstein’s inequality for the random variable $I\{A_{ir}\lor A_{jr}\}-\mathbb{E}[I\{A_{ir}\lor A_{jr}\}]$ yields that with probability at least $1-e^{-t}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}\leq\frac{4R}{\sqrt{2\pi}}+\sqrt{\frac{2t}{m}\frac{4R}{\sqrt{2\pi}}}+\frac{t}{3m}.$
Choosing $t=mR$ and taking a union bound for $i,j\in[n]$ yields that with probability at least $1-n^{2}e^{-mR}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\lor A_{jr}\}\leq 4R$ (46)
holds for all $i,j\in[n]$.
Plugging (46) into (45) leads to the conclusion. ∎
6.2. Proof of Lemma 3
Proof.
First, we can decompose $u_{i}(k+1)-u_{i}(k)$ as follows.
$u_{i}(k+1)-u_{i}(k)$ $=u_{i}(k+1)-u_{i}(k)-\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle+\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$
$:=I_{1}^{i}(k)+I_{2}^{i}(k).$
For $I_{2}^{i}(k)$, we have
$I_{2}^{i}(k)$ $=\langle\frac{\partial u_{i}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$
$=\sum\limits_{r=1}^{m}\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\rangle$
$=\sum\limits_{r=1}^{m}\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},-\eta\frac{\partial L(k)}{\partial\bm{w}_{r}}\rangle$
$=\sum\limits_{r=1}^{m}(-\eta)\sum\limits_{j=1}^{n}(u_{j}-y_{j})\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},\frac{\partial u_{j}(k)}{\partial\bm{w}_{r}}\rangle$
$=\eta\sum\limits_{j=1}^{n}(\sum\limits_{r=1}^{m}\langle\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}},\frac{\partial u_{j}(k)}{\partial\bm{w}_{r}}\rangle)(y_{j}-u_{j})$
$=\eta[\bm{H}(k)]_{i}(y-u(k)),$
where $[\bm{H}(k)]_{i}$ denotes the $i$-row of the matrix $\bm{H}(k)$.
Thus,
$u_{i}(k+1)-u_{i}(k)=\eta[\bm{H}(k)]_{i}(y-u(k))+I_{1}^{i}(k)$
and then
$(y_{i}-u_{i}(k))-(y_{i}-u_{i}(k+1))=\eta[\bm{H}(k)]_{i}(y-u(k))+I_{1}^{i}(k),$
which yields
$(\bm{y}-\bm{u}(k))-(\bm{y}-\bm{u}(k+1))=\eta\bm{H}(k)(\bm{y}-\bm{u}(k))+I_{1}(k).$
A simple transformation leads to the conclusion
$\bm{y}-\bm{u}(k+1)=(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k).$
6.3. Proof of Theorem 1
Proof.
The proof follows a similar induction to that in [10]. Our induction hypothesis is the following condition.
Condition 1.
At the $t$-th iteration, we have
$\|\bm{y}-\bm{u}(t)\|_{2}^{2}\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{t}\|\bm{y}-\bm{u}(0)\|_{2}^{2}.$ (47)
Suppose that Condition 1 holds for $t=0,\cdots,k$, then Corollary 4.1 in [10] implies that for every $r\in[m]$ and $t=0,\cdots,k$
$\|\bm{w}_{r}(t+1)-\bm{w}_{r}(0)\|_{2}\leq\frac{4\sqrt{n}\|\bm{y}-\bm{u}(0)\|_{2}}{\sqrt{m}\lambda_{0}}:=R^{{}^{\prime}}.$ (48)
In the following, we aim to show that Condition 1 also holds for $t=k+1$, thus the conclusion of Theorem 1 holds directly.
Let $S_{i}=\{r\in[m]:I\{A_{ir}\}=0\}$ and $S_{i}^{\perp}=[m]\textbackslash S_{i}$.
Recall that
$\bm{y}-\bm{u}(k+1)=(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k).$
As Condition 1 holds for $t=0,\cdots,k$, from (48), we have that for any $r\in[m]$,
$\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}.$
By taking $R=\frac{\lambda_{0}}{32n}$ in Lemma 2 and $R^{{}^{\prime}}\leq R$, we have that
$\|\bm{H}(k)-\bm{H}(0)\|_{2}\leq\frac{\lambda_{0}}{4},$
which implies that $\lambda_{min}(\bm{H}(k))\geq\lambda_{min}(\bm{H}(0))-\frac{\lambda_{0}}{4}\geq\frac{\lambda_{0}}{2}$ and
$\|\bm{H}(k)\|_{2}\leq\|\bm{H}(0)\|_{2}+\frac{\lambda_{0}}{4}\leq\|\bm{H}^{\infty}\|_{2}+\frac{\lambda_{0}}{2}\leq\frac{3}{2}\|\bm{H}^{\infty}\|_{2}.$
Thus, when $\eta\leq\frac{2}{3\|H^{\infty}\|_{2}}$, $\bm{I}-\eta\bm{H}(k)$ is positive definite and meanwhile $\|\bm{I}-\eta\bm{H}(k)\|_{2}\leq 1-\frac{\eta\lambda_{0}}{2}$. Combining with the recursion formula (10) yields that
$\|\bm{y}-\bm{u}(k+1)\|_{2}^{2}$ $=\|(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))-\bm{I}_{1}(k)\|_{2}^{2}$ (49)
$=\|(I-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}-2\langle(\bm{I}-\eta\bm{H}(k))(\bm{y}-\bm{u}(k)),\bm{I}_{1}(k)\rangle$
$\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\|\bm{y}-\bm{u}(k)\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\|\bm{I}_{1}(k)\|_{2}\|(I-\eta\bm{H}(k))(\bm{y}-\bm{u}(k))\|_{2}$
$\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\|\bm{y}-\bm{u}(k)\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2(1-\frac{\eta\lambda_{0}}{2})\|\bm{I}_{1}(k)\|_{2}\|\bm{y}-\bm{u}(k)\|_{2}.$
Then, it remains only to bound the term $\|\bm{I}_{1}(k)\|_{2}$.
Recall that for $i\in[n]$,
$I_{1}^{i}(k)=\sum\limits_{r=1}^{m}\frac{a_{r}}{\sqrt{m}}\left[\sigma(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})-\sigma(\bm{w}_{r}(k)^{T}\bm{x}_{i})-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}\right].$
Note that for $r\in S_{i}$, we can deduce that $I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{i}\geq 0\}=I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$, as $\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}\leq R$ and $\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}\leq R$. Thus, for $r\in S_{i}$, we have
$\sigma(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})-\sigma(\bm{w}_{r}(k)^{T}\bm{x}_{i})-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$
$=[(\bm{w}_{r}(k+1)^{T}\bm{x}i)I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{i}\geq 0\}-(\bm{w}_{r}(k)^{T}\bm{x}_{i})I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}]$
$\quad-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$
$=[(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}-(\bm{w}_{r}(k)^{T}\bm{x}_{i})I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}]$
$\quad-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}$
$=0.$
It implies that
$I_{1}^{i}(k)=\sum\limits_{r\in S_{i}^{\perp}}\frac{a_{r}}{\sqrt{m}}\left[\sigma(\bm{w}_{r}(k+1)^{T}\bm{x}_{i})-\sigma(\bm{w}_{r}(k)^{T}\bm{x}_{i})-(\bm{w}_{r}(k+1)-\bm{w}_{r}(k))^{T}\bm{x}_{i}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}\right].$
Therefore,
$|I_{1}^{i}(k)|$ $\leq\frac{1}{\sqrt{m}}\sum\limits_{r\in S_{i}^{\perp}}2\sqrt{2}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}$ (50)
$=\frac{2\sqrt{2}}{\sqrt{m}}\sum\limits_{r\in S_{i}^{\perp}}\|-\eta\frac{\partial L(k)}{\partial\bm{w}_{r}}\|_{2}$
$\leq\frac{4}{\sqrt{m}}\sum\limits_{r\in S_{i}^{\perp}}\frac{\eta\sqrt{n}\|\bm{y}-\bm{u}(k)\|_{2}}{\sqrt{m}}$
$=\frac{4\eta\sqrt{n}\|\bm{y}-\bm{u}(k)\|_{2}}{m}\sum\limits_{r=1}^{m}I\{r\in S_{i}^{\perp}\},$
where the second inequality follows from that
$\frac{\partial L(k)}{\partial\bm{w}_{r}}$ $=\sum\limits_{i=1}^{n}(u_{i}(k)-y_{i})\frac{\partial u_{i}(k)}{\partial\bm{w}_{r}}$
$=\sum\limits_{i=1}^{n}(u_{i}(k)-y_{i})\frac{a_{r}}{\sqrt{m}}I\{\bm{w}_{r}(k)^{T}\bm{x}_{i}\geq 0\}\bm{x}_{i}$
and then
$\left\|\frac{\partial L(k)}{\partial\bm{w}_{r}}\right\|_{2}\leq\frac{\sqrt{n}\|\bm{y}-\bm{u}(k)\|_{2}}{\sqrt{m}},$
as $\|\bm{x}_{i}\|_{2}\leq\sqrt{2}$.
Since $S_{i}=\{r\in[m]:I\{A_{ir}\}=0\}$, it follows that $I\{r\in S_{i}^{\perp}\}=I\{A_{ir}\}$, then applying Berstein’s inequality yields that with probability at least $1-ne^{-mR}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{i}^{\perp}\}\leq 4R,$ (51)
holds for all $i\in[n]$.
Thus from (50) and (51), we have
$\|I_{1}(k)\|_{2}$ $=\sqrt{\sum\limits_{i=1}^{n}|I_{1}^{i}(k)|^{2}}$ (52)
$\leq 4\eta\sqrt{n}\|y-u(k)\|_{2}\sqrt{\sum\limits_{i=1}^{n}(\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{i}^{\perp}\})^{2}}$
$\leq 8\eta nR\|y-u(k)\|_{2}$
$=\frac{\eta\lambda_{0}}{8}\|y-u(k)\|_{2},$
where the last inequality is due to $R=\frac{\lambda_{0}}{64n}$.
Plugging (52) into (49) yields that
$\|y-u(k+1)\|_{2}^{2}$ $\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\|\bm{y}-\bm{u}(k)\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2(1-\frac{\eta\lambda_{0}}{2})\|\bm{I}_{1}(k)\|_{2}\|\bm{y}-\bm{u}(k)\|_{2}$
$\leq\left[(1-\frac{\eta\lambda_{0}}{2})^{2}+\frac{\eta^{2}\lambda_{0}^{2}}{64}+2(1-\frac{\eta\lambda_{0}}{2})\frac{\eta\lambda_{0}}{8}\right]\|\bm{y}-\bm{u}(k)\|_{2}^{2}$
$\leq\left[1-\eta\lambda_{0}+\frac{\eta^{2}\lambda_{0}^{2}}{4}+\frac{\eta^{2}\lambda_{0}^{2}}{64}+\frac{\eta\lambda_{0}}{4}\right]\|\bm{y}-\bm{u}(k)\|_{2}^{2}$
$\leq(1-\frac{\eta\lambda_{0}}{2})\|\bm{y}-\bm{u}(k)\|_{2}^{2},$
where the last inequality follows from the fact that
$\eta\leq\frac{2}{3\|\bm{H}^{\infty}\|_{2}}\leq\frac{2}{3\lambda_{0}},$
which imples that
$\frac{\eta^{2}\lambda_{0}^{2}}{4}+\frac{\eta^{2}\lambda_{0}^{2}}{64}\leq\frac{\eta\lambda_{0}}{4}.$
Finally, we need consider the requirement for $m$. First, Lemma 1 shows that $m=\Omega\left(\frac{n^{2}}{\lambda_{0}^{2}}\log\left(\frac{n}{\delta}\right)\right)$. Second, from $R^{{}^{\prime}}\leq R=\frac{\lambda_{0}}{32n}$ and $R^{{}^{\prime}}=\frac{4\sqrt{n}\|\bm{y}-\bm{u}(0)\|_{2}}{\sqrt{m}\lambda_{0}}$, we know
$m=\Omega\left(\frac{n^{3}\|\bm{y}-\bm{u}(0)\|_{2}^{2}}{\lambda_{0}^{4}}\right).$
Then from the estimation in Lemma 10 for the initial prediction $\|\bm{y}-\bm{u}(0)\|_{2}$, we obtain that
$m=\Omega\left(\frac{n^{4}}{\lambda_{0}^{4}}\log\left(\frac{n}{\delta}\right)\right).$
7. Proof of Section 3
Before the proofs, we first recall that
$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{mn_{1}}}\left(\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})w_{r0}\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}\end{pmatrix}\right)$ (53)
and
$\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{mn_{2}}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j}.$ (54)
7.1. Proof of Lemma 4
Proof.
In the following, we aim to bound $\|\bm{H}(0)-\bm{H}^{\infty}\|_{F}$, as $\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}\leq\|\bm{H}(0)-\bm{H}^{\infty}\|_{F}$. Note that the entries of $\bm{H}(0)-\bm{H}^{\infty}$ have three forms as follows.
$\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle,$ (55)
$\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle$ (56)
and
$\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle.$ (57)
For the first form (55), to simplify the analysis, we let
$\bm{Z}_{r}(i)$ $=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{i})w_{r0}(0)\bm{x}_{i}+\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{i})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}$
$\quad-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\bm{x}_{p}-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{i})\begin{pmatrix}0\\ \bm{w}_{r1}(0)\end{pmatrix}$
and
$X_{r}(ij)=\langle\bm{Z}_{r}(i),\bm{Z}_{r}(j)\rangle,$
then
$\sum\limits_{r=1}^{m}\langle\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle=\frac{1}{n_{1}m}\sum\limits_{r=1}^{m}(X_{r}(ij)-\mathbb{E}X_{r}(ij)).$
Note that $|X_{r}(ij)|\lesssim 1+\|\bm{w}_{r}(0)\|_{2}^{4}$, thus
$\|X_{r}(ij)\|_{\psi_{\frac{1}{2}}}\lesssim 1+\left\|\|\bm{w}_{r}(0)\|_{2}^{4}\right\|_{\psi_{\frac{1}{2}}}\lesssim 1+\left\|\|\bm{w}_{r}(0)\|_{2}^{2}\right\|_{\psi_{1}}^{2}\lesssim d^{2}.$
For the centered random variable, the property of $\psi_{\frac{1}{2}}$ quasi-norm implies that
$\|X_{r}(ij)-\mathbb{E}[X_{r}(ij)]\|_{\psi_{\frac{1}{2}}}\lesssim\|X_{r}(ij)\|_{\psi_{\frac{1}{2}}}+\|\mathbb{E}[X_{r}(ij)]\|_{\psi_{\frac{1}{2}}}\lesssim d^{2}.$
Therefore, applying Lemma 8 yields that with probability at least $1-\delta$,
$|\sum\limits_{r=1}^{m}\frac{1}{m}(X_{r}(ij)-\mathbb{E}[X_{r}(ij)])|\lesssim\frac{d^{2}}{\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{m}(\log(\frac{1}{\delta}))^{2},$
which means
$|\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial s_{j}(\bm{w})}{\partial\bm{w}w_{r}}\rangle|\lesssim\frac{d^{2}}{n_{1}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{n_{1}m}(\log(\frac{1}{\delta}))^{2}.$ (58)
Similarly, for the second form (56) and third form (57), we can deduce that
$\left\|\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}$
and
$\left\|\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle\right\|_{\psi_{\frac{1}{2}}}\lesssim\frac{d^{2}}{n_{2}m}.$
Thus applying Lemma 8 yields that with probability at least $1-\delta$,
$|\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle|\lesssim\frac{d^{2}}{\sqrt{n_{1}n_{2}}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{\sqrt{n_{1}n_{2}}m}\log(\frac{1}{\delta})$ (59)
and with probability at least $1-\delta$,
$|\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle-\mathbb{E}_{\bm{w}}\sum\limits_{r=1}^{m}\langle\frac{\partial h_{i}(\bm{w})}{\partial\bm{w}_{r}},\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}\rangle|\lesssim\frac{d^{2}}{n_{2}\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{n_{2}m}\log(\frac{1}{\delta}).$ (60)
Combining (58), (59) and (60), we can deduce that with probability at least $1-\delta$,
$\|\bm{H}(0)-\bm{H}^{\infty}\|_{2}^{2}$
$\leq\|\bm{H}(0)-\bm{H}^{\infty}\|_{F}^{2}$
$\lesssim\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta})+\frac{d^{4}}{m^{2}}(\log(\frac{n_{1}+n_{2}}{\delta}))^{4}$
$\lesssim\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta}).$
Thus when $\sqrt{\frac{d^{4}}{m}\log(\frac{n_{1}+n_{2}}{\delta})}\lesssim\frac{\lambda_{0}}{4}$, i.e.,
$m=\Omega(\frac{d^{4}}{\lambda_{0}^{2}}\log(\frac{n_{1}+n_{2}}{\delta})),$
we have $\lambda_{min}(\bm{H}(0))\geq\frac{3}{4}\lambda_{0}$.
7.2. Proof of Lemma 5
Proof.
We first reformulate the term $\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}$ in (53) as follows.
$\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}w_{r0}(k)x_{p0}\\ w_{r0}(k)\bm{x}_{p1}-2\bm{w}_{r1}(k)\end{pmatrix}+\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\bm{x}_{p}.$
Similar to Lemma 4, it suffices to bound $\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}$, which can in turn allows us to bound each entry of $\bm{H}(\bm{w})-\bm{H}(0)$.
For $i\in[n_{1}]$ and $j\in[n_{1}]$, we have
$H_{ij}(\bm{w})$ $=\sum\limits_{r=1}^{m}\langle\frac{\partial s_{i}}{\partial\bm{w}_{r}},\frac{\partial s_{j}}{\partial\bm{w}_{r}}\rangle$
$=\frac{1}{n_{1}m}\sum\limits_{r=1}^{m}\left\langle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\begin{pmatrix}w_{r0}x_{i0}\\ w_{r0}\bm{x}_{i1}-2\bm{w}_{r1}\end{pmatrix}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{i},\right.$
$\qquad\left.\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\begin{pmatrix}w_{r0}x_{j0}\\ w_{r0}\bm{x}_{j1}-2\bm{w}_{r1}\end{pmatrix}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{j}\right\rangle$
After expanding the inner product term, we can find that although it has nine terms, it only consists of six classes. For simplicity, we use the following six symbols to represent the corresponding classes.
$\sigma^{{}^{\prime\prime}}\sigma^{{}^{\prime\prime}},\sigma^{{}^{\prime\prime}}\sigma^{{}^{\prime}},\sigma^{{}^{\prime}}\sigma^{{}^{\prime}},\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}},\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime}},\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime\prime}}.$
For instance, $\sigma^{{}^{\prime\prime}}\sigma^{{}^{\prime}}$ represents
$\left\langle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\begin{pmatrix}w_{r0}x_{i0}\\ w_{r0}\bm{x}_{i1}-2\bm{w}_{r1}\end{pmatrix},\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix}\right\rangle,\left\langle\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\begin{pmatrix}1\\ \bm{0}_{d+1}\end{pmatrix},\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\begin{pmatrix}w_{r0}x_{j0}\\ w_{r0}\bm{x}_{j1}-2\bm{w}_{r1}\end{pmatrix}\right\rangle.$
In fact, when bounding the corresponding terms for $H_{ij}(\bm{w})-H_{ij}(0)$, the first four classes can be grouped into one category. They are of the form $f_{1}(\bm{w})f_{2}(\bm{w})f_{3}(\bm{w})f_{4}(\bm{w})$, where for each $1\leq i\leq 4$, $f_{i}(\bm{w})$ is Lipschitz continuous with respect to $\|\cdot\|_{2}$ and $|f_{i}(\bm{w})|\lesssim\|\bm{w}\|_{2}$ (Note that $\sigma^{{}^{\prime}}(\cdot)=(\sigma^{{}^{\prime\prime}}(\cdot))^{2}$). On the other hand, when $\|\bm{w}_{1}-\bm{w}_{2}\|_{2}\leq R\leq 1$, we can deduce that
$|f_{1}(\bm{w}_{1})f_{2}(\bm{w}_{1})f_{3}(\bm{w}_{1})f_{4}(\bm{w}_{1})-f_{1}(\bm{w}_{2})f_{2}(\bm{w}_{2})f_{3}(\bm{w}_{2})f_{4}(\bm{w}_{2})|\lesssim R(\|\bm{w}_{1}\|_{2}^{3}+1).$
Thus, for the terms in $H_{ij}(w)-H_{ij}(0)$ that belong to the first four classes, we can deduce that they are less than $CR(\|\bm{w}_{r}(0)\|_{2}^{3}+1)$, where $C$ is a universal constant.
For the classes $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime}}$, they are both involving $\sigma^{{}^{\prime\prime\prime}}$ that is not Lipschitz continuous. To make it precise, we write the class $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}}$ explicitly as follows.
$\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\|\bm{w}_{r1}\|_{2}^{2}\begin{pmatrix}w_{r0}x_{i0}\\ w_{r0}\bm{x}_{i1}-2\bm{w}_{r1}\end{pmatrix}^{T}\bm{x}_{j}.$
Note that when $\|\bm{w}_{r}-\bm{w}_{r}(0)\|_{2}<R$, we have that
$|\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{j})|=|I\{\bm{w}_{r}^{T}\bm{x}_{j}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{j}\geq 0\}|\leq I\{A_{jr}\},$
where the definition of $A_{jr}$ is same as (44) for the regression problems.
Thus, we can deduce that for the terms in $H_{ij}(w)-H_{ij}(0)$ that belong to the classes $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime}}$, they are less than
$C\left[(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{3}+R(\|\bm{w}_{r}(0)\|_{2}^{3}+1)\right],$
where $C$ is a universal constant.
Similarly, for the last class $\sigma^{{}^{\prime\prime\prime}}\sigma^{{}^{\prime\prime\prime}}$ that are of the form
$\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\|\bm{w}_{r1}\|_{2}^{4}\bm{x}_{i}^{T}\bm{x}_{j},$
we can deduce that
$|\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{i})\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{j})\|\bm{w}_{r1}\|_{2}^{4}\bm{x}_{i}^{T}\bm{x}_{j}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{i})\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{j})\|\bm{w}_{r1}(0)\|_{2}^{4}\bm{x}_{i}^{T}\bm{x}_{j}|$
$\lesssim I\{A_{ir}\lor A_{jr}\}\|\bm{w}_{r}(0)\|_{2}^{4}+R\|\bm{w}_{r}(0)\|_{2}^{3}+R.$
Combining the upper bounds for the six classes, we have that
$|H_{ij}(\bm{w})-H_{ij}(0)|$ $\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{3})+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{4}+R\right]$ (61)
$\lesssim\frac{1}{n_{1}}\left[\frac{1}{m}(R\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4})+\frac{1}{m}\sum\limits_{r=1}^{m}(I\{A_{ir}\}+I\{A_{jr}\})\|\bm{w}_{r}(0)\|_{2}^{4}+R\right],$
where the last inequality follows from that $\|\bm{w}_{r}(0)\|_{2}^{3}\lesssim\|\bm{w}_{r}(0)\|_{2}^{4}+1$ due to Young’s inequality for products.
Now, we focus on the term $\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|\bm{w}_{r}(0)\|_{2}^{4}$.
Since
$P(|w_{ri}(0)|^{2}\geq 2\log(\frac{2}{\delta}))\leq\delta$
and then
$P(\|w_{r}(0)\|_{2}^{2}\geq 2(d+2)\log(\frac{2(d+2)}{\delta}))\leq\delta,$
we obtain that
$P(\exists r\in[m],\|w_{r}(0)\|_{2}^{2}\geq 2(d+2)\log\frac{2m(d+2)}{\delta})\leq\delta.$ (62)
Let $M=2(d+2)\log(\frac{2m(d+2)}{\delta})$, then
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}$
$=\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}\leq M\}+\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}>M\}$
$\leq\frac{M^{2}}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|w_{r}(0)\|_{2}^{4}I\{\|w_{r}(0)\|_{2}^{2}>M\}.$
Applying Bernstein’s inequality for the first term yields that with probability at least $1-e^{-mR}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{ir}\}\leq 4R.$
Moreover, with probability at least $1-\delta$, $I\{\|w_{r}(0)\|_{2}^{2}>M\}=0$ holds for all $r\in[m]$.
Thus from (61), with probability at least $1-\delta-n_{1}e^{-mR}$, we have that for any $i\in[n_{1}]$ and $j\in[n_{1}]$,
$|H_{ij}(\bm{w})-H_{ij}(0)|$ $\lesssim\frac{1}{n_{1}}\left[RM^{2}+RM^{2}+R\right]$
$\lesssim\frac{1}{n_{1}}M^{2}R.$
For $i\in[n_{1}],j\in[n_{1}+2,n_{2}]$ and $i\in[n_{1}+1,n_{2}],j\in[n_{2}]$, from the form of $\frac{\partial h_{j}}{\partial\bm{w}_{r}}$, i.e.,
$\frac{\partial h_{j}}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{n_{2}m}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j},$
we can obtain similar results for the terms $\langle\frac{\partial s_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\rangle$ and $\langle\frac{\partial h_{i}}{\partial\bm{w}},\frac{\partial h_{j}}{\partial\bm{w}}\rangle$.
With all results above, we have that with probability at least $1-\delta-(n_{1}+n_{2})e^{-mR}$,
$\|\bm{H}(\bm{w})-\bm{H}(0)\|_{F}\lesssim M^{2}R.$
7.3. Proof of Lemma 6
Proof.
Similar to the proof of Lemma 3, we have
$s_{p}(k+1)-s_{p}(k)$ $=\left[s_{p}(k+1)-s_{p}(k)-\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle\right]+\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (63)
$:=I_{1}^{p}(k)+I_{2}^{p}(k).$
For the second term $I_{2}^{p}(k)$, we have that
$I_{2}^{p}(k)$ $=\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (64)
$=\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},-\eta\frac{\partial L(k)}{\partial\bm{w}}\rangle$
$=-\sum\limits_{r=1}^{m}\eta\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\frac{\partial L(k)}{\partial\bm{w}_{r}}\rangle$
$=-\sum\limits_{r=1}^{m}\eta\langle\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}},\sum\limits_{t=1}^{n_{1}}s_{t}(k)\frac{\partial s_{t}(k)}{\partial\bm{w}_{r}}+\lambda\sum\limits_{j=1}^{n_{2}}h_{j}(k)\frac{\partial h_{j}(k)}{\partial\bm{w}_{r}}\rangle$
$=-\eta[\bm{H}(k)]_{p}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},$
where $[\bm{H}(k)]_{p}$ denotes the $p$-row of $\bm{H}(k)$.
Similarly, for $h(k)$, we have
$h_{j}(k+1)-h_{j}(k)$ $=\left[h_{j}(k+1)-h_{j}(k)-\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle\right]+\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle$ (65)
$:=I_{1}^{n_{1}+j}(k)+I_{2}^{n_{1}+j}(k)$
and
$I_{2}^{n_{1}+j}(k)=-\eta[\bm{H}(k)]_{n_{1}+j}\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}.$ (66)
Combining (63), (64), (65) and (66) yields that
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}-\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}$ $=\bm{I}_{1}(k)+\bm{I}_{2}(k)$
$=\bm{I}_{1}(k)-\eta\bm{H}(k)\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}.$
A simple transformation directly leads to
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k).$
7.4. Proof of Theorem 2
Proof.
The proof strategy is similar to that for Theorem 1. Our induction hypothesis is the following boundedness of the hidden weights and convergence rate of the empirical loss.
Condition 2.
At the $t$-th iteration, we have that for each $r\in[m]$, $\|\bm{w}_{r}(t)\|_{2}\leq B$ and
$L(t)\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{t}L(0),$ (67)
where $B=\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}+1$.
From (62), we know that with probability at least $1-\delta$, $\|\bm{w}_{r}(0)\|_{2}\leq\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}$ holds for any $r\in[m]$. Thus, if we can prove that $\bm{w}_{r}(t)$ is closed enough to $\bm{w}_{r}(0)$, then $\|\bm{w}_{r}(t)\|_{2}\leq B$.
Corollary 2 (Lemma 4.1 in [19]).
If Condition 2 holds for $t=0,\cdots,k$, then we have for every $r\in[m]$,
$\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}:=R^{{}^{\prime}},$ (68)
where $C$ is a universal constant.
Corollary 2 implies that when $m$ is large enough, we have $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq 1$ and then $\|\bm{w}_{r}(k+1)\|_{2}\leq B$. Thus, in induction, we only need to prove that (67) also holds for $t=k+1$, which relies on the recursion formula (26).
Recall that
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k).$
From Corollary 2 and Lemma 5, taking $CM^{2}R<\frac{\lambda_{0}}{4}$ in (24) and $R^{{}^{\prime}}\leq R$ in (68) yields that $\lambda_{min}(\bm{H}(k))\geq\frac{\lambda_{0}}{2}$ and
$\|\bm{H}(k)\|_{2}\leq\|\bm{H}(0)\|_{2}+\frac{\lambda_{0}}{4}\leq\|\bm{H}^{\infty}\|_{2}+\frac{\lambda_{0}}{2}\leq\frac{3}{2}\|\bm{H}^{\infty}\|_{2}.$
Therefore, if we take $\eta\leq\frac{2}{3}\frac{1}{\|\bm{H}^{\infty}\|_{2}}$, then $\bm{I}-\eta\bm{H}(k)$ is positive definite and $\|I-\eta\bm{H}(k)\|_{2}\leq 1-\frac{\eta\lambda_{0}}{2}$.
Combining these facts with the recursion formula, we have that
$\left\|\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}\right\|_{2}^{2}$ (69)
$=\left\|(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left\langle(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},\bm{I}_{1}(k)\right\rangle$
$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}.$
Thus, it remains to bound $\|\bm{I}_{1}(k)\|_{2}$.
For $\bm{I}_{1}(k)$, recall that $\bm{I}_{1}(k)=(I_{1}^{1}(k),\cdots,I_{1}^{n_{1}}(k),I_{1}^{n_{1}+1}(k),\cdots,I_{1}^{n_{1}+n_{2}}(k))^{T}\in\mathbb{R}^{n_{1}+n_{2}}$ and for $p\in[n_{1}]$,
$I_{1}^{p}(k)=s_{p}(k+1)-s_{p}(k)-\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle,$
for $j\in[n_{2}]$,
$I_{1}^{n_{1}+j}(k)=h_{j}(k+1)-h_{j}(k)-\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle.$
Recall that
$s_{p}(k)=\frac{1}{\sqrt{n_{1}}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}x_{p})w_{r0}(k)-a_{r}\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}x_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}-f(x_{p})\right)$
and
$\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+2}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\bm{x}_{p}$
$\quad-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}(k)\end{pmatrix}].$
Define $\chi_{pr}^{1}(k):=\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)$ and $\chi_{pr}^{2}(k):=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$, i.e., $\chi_{pr}^{1}(k)$ and $\chi_{pr}^{2}(k)$ are related to the operators $\frac{\partial u}{\partial t}$ and $\Delta u$ respectively.
Then define
$\hat{\chi}_{pr}^{1}(k)=\chi_{pr}^{1}(k+1)-\chi_{pr}^{1}(k)-\langle\frac{\partial\chi_{pr}^{1}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\rangle$
and
$\hat{\chi}_{pr}^{2}(k)=\chi_{pr}^{2}(k+1)-\chi_{pr}^{2}(k)-\langle\frac{\partial\chi_{pr}^{2}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\rangle.$
At this time, we have
$I_{1}^{p}(k)=\frac{1}{\sqrt{n_{1}m}}\sum\limits_{r=1}^{m}a_{r}\left[\hat{\chi}_{pr}^{1}(k)-\hat{\chi}_{pr}^{2}(k)\right].$
The purpose of defining them in this way is to enable us to handle the terms related to the operators $\frac{\partial u}{\partial t}$ and $\Delta u$ separately.
Similar to that for regression problems with ReLU activation function, we first define some events. For $p\in[n_{1}]$, define
$A_{p,r}=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}(0)\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{p}\geq 0\}\neq I\{\bm{w}_{r}(0)^{T}\bm{x}_{p}\geq 0\}\}$
and let $S_{p}^{\perp}=[n_{1}]\textbackslash S_{p}$.
In the following, we are going to show that $|\hat{\chi}_{pr}^{1}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$ for every $r\in[m]$ and $|\hat{\chi}_{pr}^{2}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$ for $r\in S_{p}$, $|\hat{\chi}_{pr}^{2}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2})$ for $r\in S_{p}^{\perp}$. Thus, we can prove that $\|\bm{I}_{1}\|_{2}=\mathcal{O}\left(\frac{\sqrt{L(k)}}{\sqrt{m}}\right)$. Then combining with (69) leads to a milder requirement for $\eta$.
For $\hat{\chi}_{pr}^{1}(k)$, we have that
$\hat{\chi}_{pr}^{1}(k)$ $=\sigma^{{}^{\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})w_{r0}(k+1)-\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)$
$\quad-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)-(w_{r0}(k+1)-w_{r0}(k))\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})$
$=(\sigma^{{}^{\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p}))w_{r0}(k+1)-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k).$
From the mean value theorem, we can deduce that there exists $\zeta(k)\in\mathbb{R}$ such that
$\sigma^{{}^{\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})=\sigma^{{}^{\prime\prime}}(\zeta(k))\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle$
and
$|\sigma^{{}^{\prime\prime}}(\zeta(k))-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})|$ $\leq|\zeta(k)-\bm{w}_{r}(k)^{T}\bm{x}_{p}|$
$\leq\sqrt{2}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}.$
Then, for $\hat{\chi}_{pr}^{1}(k)$, we can rewrite it as follows.
$\hat{\chi}_{pr}^{1}(k)$ $=\sigma^{{}^{\prime\prime}}(\zeta(k))\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle w_{r0}(k+1)-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)$
$=\left[\left(\sigma^{{}^{\prime\prime}}(\zeta(k))-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\right)\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle w_{r0}(k+1)\right]$
$\quad+\left[\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})(w_{r0}(k+1)-w_{r0}(k))\right].$
It implies that
$|\hat{\chi}_{pr}^{1}(k)|\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}.$
For $\hat{\chi}_{pr}^{2}(k)$, we write it as follows explicitly.
$\hat{\chi}_{pr}^{2}(k)$ $=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k+1)\|_{2}^{2}-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$ (70)
$\quad-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$
$\quad-2\langle\bm{w}_{r1}(k+1)-\bm{w}_{r1}(k),\bm{w}_{r1}(k)\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p}).$
Note that for the term $\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{w}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$, we can rewrite it as follows.
$\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$ (71)
$=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)-\bm{w}_{r1}(k+1)+\bm{w}_{r1}(k+1)\|_{2}^{2}$
$=\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})[\|\bm{w}_{r1}(k)-\bm{w}_{r1}(k+1)\|_{2}^{2}+\|\bm{w}_{r1}(k+1)\|_{2}^{2}-2\langle\bm{w}_{r1}(k+1)-\bm{w}_{r1}(k),\bm{w}_{r1}(k+1)\rangle],$
where the first term $\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)-\bm{w}_{r1}(k+1)\|_{2}^{2}=\mathcal{O}(B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$.
plugging (71) into (70) yields that
$\hat{\chi}_{pr}^{2}(k)$ $=[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})]\|\bm{w}_{r1}(k+1)\|_{2}^{2}$ (72)
$\quad-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}$
$\quad+2\langle\bm{w}_{r1}(k+1)-\bm{w}_{r1}(k),\bm{w}_{r1}(k+1)-\bm{w}_{r1}(k)\rangle\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})+\mathcal{O}(B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$
$=[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})]\|\bm{w}_{r1}(k+1)\|_{2}^{2}$
$\quad+\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})(\|\bm{w}_{r1}(k+1)\|_{2}^{2}-\|\bm{w}_{r1}(k)\|_{2}^{2})$
$\quad+\mathcal{O}(B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$
$=\left[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\right]\|\bm{w}_{r1}(k+1)\|_{2}^{2}$
$\quad+\mathcal{O}(B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}).$
Thus, we only need to consider the term
$\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p}).$
For $r\in S_{p}$, as $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq R,\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R$, we have that $I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{p}\geq 0\}=I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}$, which yields that
$\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})$ (73)
$=[(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{p}\geq 0\}-(\bm{w}_{r}(k)^{T}\bm{x}_{p})I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}]$
$\quad-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}$
$=[(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}-(\bm{w}_{r}(k)^{T}\bm{x}_{p})I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}]$
$\quad-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}$
$=0.$
For $r\in S_{p}^{\perp}$, the Lipschitz continuity of $\sigma^{{}^{\prime\prime}}$ implies that
$\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k+1)^{T}\bm{x}_{p})-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{x}_{p}\rangle\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}).$ (74)
Combining (72), (73) and (74), we can deduce that for $r\in S_{p}$,
$|\hat{\chi}_{pr}^{2}(k)|\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2},$
for $r\in S_{p}^{\perp}$,
$|\hat{\chi}_{pr}^{2}(k)|\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}+B^{2}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}.$
With the estimations for $\hat{\chi}_{pr}^{1}(k)$ and $\hat{\chi}_{pr}^{2}(k)$, we have
$|I_{1}^{p}(k)|$ $\leq\frac{1}{\sqrt{n_{1}m}}\sum\limits_{r=1}^{m}(|\hat{\chi}_{pr}^{1}(k)|+|\hat{\chi}_{pr}^{2}(k)|)$ (75)
$\lesssim\frac{1}{\sqrt{n_{1}m}}\sum\limits_{r=1}^{m}B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}+\frac{1}{\sqrt{n_{1}m}}\sum\limits_{r\in S_{p}^{\perp}}B^{2}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}.$
For $j\in[n_{2}]$, we consider $I_{1}^{n_{1}+j}(k)$, which can be written as follows.
$I_{1}^{n_{1}+j}(k)$ $=h_{j}(k+1)-h_{j}(k)-\langle\bm{w}(k+1)-\bm{w}(k),\frac{\partial h_{j}(k)}{\partial\bm{w}}\rangle$
$=\sum\limits_{r=1}^{m}\frac{a_{r}}{\sqrt{n_{2}m}}\left[\sigma(\bm{w}_{r}(k+1)^{T}\bm{y}_{j})-\sigma(\bm{w}_{r}(k)^{T}\bm{y}_{j})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{y}_{j})\right].$
From the mean value theorem, we have that there exists $\zeta(k)\in\mathbb{R}$ such that
$\sigma(\bm{w}_{r}(k+1)^{T}\bm{y}_{j})-\sigma(\bm{w}_{r}(k)^{T}\bm{y}_{j})=\sigma^{{}^{\prime}}(\zeta(k))\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle$
and
$|\sigma^{{}^{\prime}}(\zeta(k))-\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{y}_{j})|$ $\leq 2B|\zeta(k)-\bm{w}_{r}(k)^{T}\bm{y}_{j}|$
$\leq 2\sqrt{2}B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}.$
Thus,
$|\sigma(\bm{w}_{r}(k+1)^{T}\bm{y}_{j})-\sigma(\bm{w}_{r}(k)^{T}\bm{y}_{j})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{y}_{j})|$
$=|\sigma^{{}^{\prime}}(\zeta(k))\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle-\sigma(\bm{w}_{r}(k)^{T}\bm{y}_{j})-\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{y}_{j})|$
$=|(\sigma^{{}^{\prime}}(\zeta(k))-\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{y}_{j}))\langle\bm{w}_{r}(k+1)-\bm{w}_{r}(k),\bm{y}_{j}\rangle|$
$\lesssim B\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}.$
Therefore, for $j\in[n_{2}]$,
$|I_{1}^{n_{1}+j}(k)|\lesssim\frac{B}{\sqrt{n_{2}m}}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}.$ (76)
From the updating rule of gradient descent, we can deduce that for every $r\in[m]$,
$\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}=\left\|-\eta\frac{\partial L(k)}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{\eta B^{2}}{\sqrt{m}}\sqrt{L(k)}.$ (77)
Plugging (77) into (75) and (76), we can deduce that
$|I_{1}^{p}(k)|$ $\lesssim\frac{B}{\sqrt{n_{1}m}}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}+\frac{B^{2}}{\sqrt{n_{1}m}}\sum\limits_{r\in S_{p}^{\perp}}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}$ (78)
$\lesssim\frac{B}{\sqrt{n_{1}m}}\sum\limits_{r=1}^{m}\frac{\eta^{2}B^{4}}{m}L(k)+\frac{B^{2}}{\sqrt{n_{1}m}}\sum\limits_{r\in S_{p}^{\perp}}\frac{\eta B^{2}}{\sqrt{m}}\sqrt{L(k)}$
$=\frac{\eta^{2}B^{5}L(k)}{\sqrt{n_{1}m}}+\frac{\eta B^{4}\sqrt{L(k)}}{\sqrt{n_{1}}}\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{p}^{\perp}\}$
$\leq\frac{\eta^{2}B^{5}\sqrt{L(0)}\sqrt{L(k)}}{\sqrt{n_{1}m}}+\frac{\eta B^{4}\sqrt{L(k)}}{\sqrt{n_{1}}}\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{p}^{\perp}\}$
and
$|I_{1}^{n_{1}+j}(k)|$ $\lesssim\frac{B}{\sqrt{n_{2}m}}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2}$ (79)
$\lesssim\frac{B}{\sqrt{n_{2}m}}\sum\limits_{r=1}^{m}\frac{\eta^{2}B^{4}}{m}L(k)$
$\leq\frac{\eta^{2}B^{5}\sqrt{L(0)}\sqrt{L(k)}}{\sqrt{n_{2}m}}.$
Recall that
$A_{pr}=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{p}\geq 0\}\neq I\{\bm{w}_{r}^{T}\bm{x}_{p}\geq 0\}\}$
and
$P(A_{p,r})\leq\frac{R}{\sqrt{2\pi}},\ S_{p}=\{r\in[m]:I\{A_{p,r}\}=0\}.$
Thus, from Bernstein’s inequality, we have that with probability at least $1-e^{-mR}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{r\in S_{p}^{\perp}\}=\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{pr}\}\lesssim 4R.$
Then the inequality holds for all $p\in[n_{1}]$ with probability at least $1-n_{1}e^{-mR}$. Thus from (78), we have that for every $p\in[n_{1}]$
$|I_{1}^{p}(k)|\lesssim\frac{\eta^{2}B^{5}\sqrt{L(0)}\sqrt{L(k)}}{\sqrt{n_{1}m}}+\frac{\eta B^{4}\sqrt{L(k)}}{\sqrt{n_{1}}}R.$ (80)
Combining (79) and (80), we have that
$\|\bm{I}_{1}(k)\|_{2}$ $=\sqrt{\sum\limits_{p=1}^{n_{1}}|I_{1}^{p}(k)|^{2}+\sum\limits_{j=1}^{n_{2}}|I_{1}^{n_{1}+j}(k)|^{2}}$
$\lesssim\frac{\eta^{2}B^{5}\sqrt{L(0)}\sqrt{L(k)}}{\sqrt{m}}+\eta B^{4}\sqrt{L(k)}R.$
Plugging it into (69) yields that
$\left\|\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}\right\|_{2}^{2}$
$\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2(1-\frac{\eta\lambda_{0}}{2})\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}$
$\leq[(1-\frac{\eta\lambda_{0}}{2})^{2}+C^{2}(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R)^{2}+2C(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R)]\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}$
$\leq(1-\frac{\eta\lambda_{0}}{2})\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2},$
where the last inequality requires that
$\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}\lesssim\eta\lambda_{0},\ \eta B^{4}R\lesssim\eta\lambda_{0}.$
Recall that we also require that $CM^{2}R<\frac{\lambda_{0}}{4}$ and
$R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}<R.$
Finally, with $R=\mathcal{O}(\frac{\lambda_{0}}{B^{4}})$ and Lemma 11 for the upper bound of $L(0)$, $m$ needs to satisfies that
$m=\Omega\left(\frac{B^{12}L(0)}{\lambda_{0}^{4}}\right)=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
8. Proof of Section 4
8.1. Proof of Lemma 7
Proof.
Recall that
$\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}x_{p})w_{r0}\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\|\bm{w}_{r1}\|_{2}^{2}\bm{x}_{p}$
$\quad-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}\end{pmatrix}]$
and
$\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}=\frac{a_{r}}{\sqrt{n_{2}m}}\sigma^{{}^{\prime}}(\bm{w}_{r}^{T}\bm{y}_{j})\bm{y}_{j}.$
(1) For $p\in[n_{1}]$, define the event
$A_{p,r}=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}(0)\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{p}\geq 0\}\neq I\{\bm{w}_{r}(0)^{T}\bm{x}_{p}\geq 0\}\}.$
Similar to (44), we can deduce that for any $p\in[n_{1}]$,
$P(A_{p,r})\leq\frac{2R}{\sqrt{2\pi}}.$
From the form of $\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}$, we can deduce that
$\left\|\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(0)}{\partial\bm{w}_{r}}\right\|_{2}$ (81)
$\lesssim\frac{1}{\sqrt{n_{1}m}}\left[R(\|\bm{w}_{r}(0)\|_{2}+1)+|I\{\bm{w}_{r}^{T}\bm{x}_{p}\geq 0\}-I\{\bm{w}_{r}(0)^{T}\bm{x}_{p}\geq 0\}|(\|\bm{w}_{r}(0)\|_{2}^{2}+1)\right]$
$\leq\frac{1}{\sqrt{n_{1}m}}\left[R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)\right],$
where the second inequality follows from the fact $\|\bm{w}-\bm{w}_{r}(0)\|_{2}<R\leq 1$ and the definition of $A_{p,r}$.
Similarly, we have that
$\left\|\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(0)}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{2}m}}R(\|\bm{w}_{r}(0)\|_{2}+1).$ (82)
Combining (81) and (82), we can deduce that
$\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}^{2}$
$\leq\|\bm{J}(\bm{w})-\bm{J}(0)\|_{F}^{2}$
$=\sum\limits_{i=1}^{n_{1}+n_{2}}\|\bm{J}_{i}(\bm{w})-\bm{J}_{i}(0)\|_{2}^{2}$
$=\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(0)}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\sum\limits_{j=1}^{n_{2}}\left\|\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(0)}{\partial\bm{w}_{r}}\right\|_{2}^{2}\right)$
$\lesssim\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\frac{1}{n_{1}m}(R(\|\bm{w}_{r}(0)\|_{2}+1)+I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{2}+1))^{2}+\sum\limits_{j=1}^{n_{2}}\frac{1}{n_{2}m}(R\|\bm{w}_{r}(0)\|_{2}+R)^{2}\right)$
$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}(\|\bm{w}_{r}(0)\|_{2}^{4}+1)$
$=\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)$
$\quad+\frac{1}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}\left(\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}\leq M\}+\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\}+1\right)$
$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{2}+1)+\frac{M^{2}}{n_{1}m}\sum\limits_{p=1}^{n_{1}}\sum\limits_{r=1}^{m}I\{A_{p,r}\}+\frac{1}{m}\sum\limits_{r=1}^{m}\|\bm{w}_{r}(0)\|_{2}^{4}I\{\|\bm{w}_{r}(0)\|_{2}^{2}>M\},$
where $M=2(d+2)\log(2m(d+2)/\delta)$. Note that from (62), we have
$P\left(\exists r\in[m],\|\bm{w}_{r}(0)\|_{2}^{2}\geq 2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)\right)\leq\delta.$
On the other hand, applying Bernstein’s inequality yields that with probability at least $1-n_{1}e^{-mR}$,
$\frac{1}{m}\sum\limits_{r=1}^{m}I\{A_{p,r}\}<4R$
holds for all $p\in[n_{1}]$.
Therefore, we have that
$\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}^{2}\lesssim MR^{2}+R^{2}+M^{2}R\lesssim M^{2}R$
holds with probability at least $1-\delta-n_{1}e^{-mR}$.
(2) Note that when $\sigma$ satisfies the assumption 5, $\sigma^{{}^{\prime}},\sigma^{{}^{\prime\prime}}$ and $\sigma^{{}^{\prime\prime\prime}}$ are all Lipschitz continuous. Thus we can obtain that
$\left\|\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(0)}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{1}m}}R(\|\bm{w}_{r}(0)\|_{2}^{2}+\|\bm{w}_{r}(0)\|_{2}+1)\lesssim\frac{1}{\sqrt{n_{1}m}}R(\|\bm{w}_{r}(0)\|_{2}^{2}+1),$ (83)
where the second inequality is from Young’s inequality.
Similarly, we have
$\left\|\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(0)}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{2}m}}R(\|\bm{w}_{r}(0)\|_{2}+1).$ (84)
Combining (83) and (84) yields that
$\|\bm{J}(\bm{w})-\bm{J}(0)\|_{2}^{2}$
$\leq\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(0)}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\sum\limits_{j=1}^{n_{2}}\left\|\frac{\partial h_{j}(\bm{w})}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(0)}{\partial\bm{w}_{r}}\right\|_{2}^{2}\right)$
$\lesssim\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\frac{1}{n_{1}m}(R\|\bm{w}_{r}(0)\|_{2}^{2}+R)^{2}+\sum\limits_{j=1}^{n_{2}}\frac{1}{n_{2}m}(R\|\bm{w}_{r}(0)\|_{2}+R)^{2}\right)$
$\lesssim\frac{R^{2}}{m}\sum\limits_{r=1}^{m}(\|\bm{w}_{r}(0)\|_{2}^{4}+1)$
$\lesssim R^{2}\left[d^{2}+\frac{d^{2}}{\sqrt{m}}\sqrt{\log(\frac{1}{\delta})}+\frac{d^{2}}{m}(\log(\frac{1}{\delta}))^{2}\right],$
where the last inequality follows from the fact that $\|\|\bm{w}_{r}(0)\|_{2}^{4}\|_{\psi_{\frac{1}{2}}}\lesssim d^{2}$ and Lemma 8. ∎
8.2. Proof of Theorem 3
We prove Theorem 3 by induction. Our induction is the following condition for the hidden weights.
Condition 3.
At the $t$-th iteration, we have $\|\bm{w}_{r}(t)\|_{2}\leq B$ and
$\|\bm{w}_{r}(t)-\bm{w}_{r}(0)\|_{2}\leq\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}:=R^{{}^{\prime}}$
for all $r\in[m]$, where $C$ is a universal constant and $B=\sqrt{2(d+2)\log\left(\frac{2m(d+2)}{\delta}\right)}+1.$
Instead of inducing on the convergence rate of the empirical loss function, as shown in Condition 2, we perform induction on the movements of the hidden weights. Because with Condition 3, we can directly derive the following convergence rate of the empirical loss function.
Corollary 3.
If Condition 3 holds for $t=0,\cdots,k$ and $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\lesssim\sqrt{\lambda_{0}}$, then
$L(t)\leq(1-\eta)^{t}L(0),$
holds for $t=0,\cdots,k$, where $R$ is the constant in Lemma 7 and $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (36) when $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CdR$ in (38) when $\sigma$ satisfies Asssumption 5.
Thanks to Corollary 3, it is sufficient to prove that Condition 3 also holds for $t=k+1$. For readability, we defer the proof of Corollary 3 to the end of this section. In the following, we are going to show that the Condition 3 also holds for $t=k+1$, thus combining Condition 3 and Corollary 3 leads to Theorem 3.
Proof of Theorem 3.
We let $R^{{}^{\prime\prime}}=CM\sqrt{R}$ in (36) when $\sigma$ is the $\text{ReLU}^{3}$ activation function and let $R^{{}^{\prime\prime}}=CdR$ in (37) when $\sigma$ satisfies Assumption 5.
First, set $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$, then from Lemma 7 we have $\|\bm{J}(t)-\bm{J}(0)\|_{2}\leq\frac{\sqrt{\lambda_{0}}}{3}$, thus $\lambda_{min}(\bm{J}(t))\geq\frac{2\sqrt{\lambda_{0}}}{3}$ and $\lambda_{min}(\bm{H}(t))\geq\frac{4\lambda_{0}}{9}$ for $t=0,\cdots,k$.
From the updating role of NGD, we have
$\bm{w}_{r}(t+1)=\bm{w}_{r}(t)-\eta\left[\bm{J}(t)^{T}\right]_{r}(\bm{H}(t))^{-1}\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix},$
where
$\left[\bm{J}(t)^{T}\right]_{r}=\left(\frac{\partial s_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial s_{n_{1}}(t)}{\partial\bm{w}_{r}},\frac{\partial h_{1}(t)}{\partial\bm{w}_{r}},\cdots,\frac{\partial h_{n_{2}}(t)}{\partial\bm{w}_{r}}\right).$
Therefore, for $t=0,\cdots,k$ and any $r\in[m]$, we have
$\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}$ (85)
$\leq\eta\|\left[\bm{J}(t)^{T}\right]_{r}\|_{2}\|\bm{H}(t)^{-1}\|_{2}\sqrt{L(t)}$
$\leq\frac{9\eta}{4\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{2}\sqrt{L(t)}$
$\leq\frac{9\eta}{4\lambda_{0}}\|\left[\bm{J}(t)^{T}\right]_{r}\|_{F}\sqrt{L(t)}$
$=\frac{9\eta}{4\lambda_{0}}\sqrt{\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\sum\limits_{j=1}^{n_{2}}\left\|\frac{\partial h_{j}(t)}{\partial\bm{w}_{r}}\right\|_{2}^{2}}\sqrt{L(t)}$
$\lesssim\frac{4\eta}{9\lambda_{0}}\sqrt{\frac{B^{4}+1}{m}}\sqrt{L(t)}$
$\lesssim\frac{\eta B^{2}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}$
$\leq\frac{\eta B^{2}}{\sqrt{m}\lambda_{0}}(1-\eta)^{t/2}\sqrt{L(0)},$
where the last inequality is due to Corollary 3.
Summing $t$ from 00 to $k$ yields that
$\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}$
$\leq\sum\limits_{t=0}^{k}\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}$
$\leq C\frac{\eta B^{2}}{\sqrt{m}\lambda_{0}}\sum\limits_{t=0}^{k}(1-\eta)^{t/2}\sqrt{L(0)}$
$\leq\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}.$
Note that although $C$ is a universal constant, it still depends on the activation function.
Now, when $R^{{}^{\prime}}\leq 1$, we can deduce that Condition 3 also holds for $t=k+1$. Thus, it remains only to derive the requirement for $m$.
Recall that we need $m$ to satisfy that $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$.
(1) When $\sigma$ is the $\text{ReLU}^{3}$ activation function, $R^{{}^{\prime\prime}}=CM\sqrt{R}\leq\frac{\sqrt{\lambda_{0}}}{3}$, wihch implies that $R\lesssim\frac{\lambda_{0}}{M^{2}}$. Then $R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R$ imples that
$m=\Omega\left(\frac{M^{4}B^{4}L(0)}{\lambda_{0}^{4}}\right).$
From Lemma 11, we can deduce that
$m=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
(2) When $\sigma$ satisfies Assumption 5, we have that
$R\lesssim\frac{\sqrt{\lambda_{0}}}{d},R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}\leq R.$
From Lemma 5 in [23], we know
$L(0)\lesssim d^{2}\log\left(\frac{n_{1}+n_{2}}{\delta}\right).$
Thus, we can deduce that
$m=\Omega\left(\frac{d^{6}}{\lambda_{0}^{3}}\log^{2}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
Proof of Corollary 3.
Similar as before, when $R^{{}^{\prime}}\leq R$ and $R^{{}^{\prime\prime}}\leq\frac{\sqrt{\lambda_{0}}}{3}$ , we have $\lambda_{min}(\bm{J}(t))\geq\frac{2\sqrt{\lambda_{0}}}{3}$ and then $\lambda_{min}(\bm{H}(t))\geq\frac{4}{9}\lambda_{0}$ for $t=0,\cdots,k$.
Let $\bm{u}(t)=\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}$, then
$\bm{u}(t+1)-\bm{u}(t)$ (86)
$=\bm{u}\left(\bm{w}(t)-\eta\bm{J}(t)^{T}\bm{H}(t)^{-1}\bm{u}(\bm{w}(t))\right)-\bm{u}(\bm{w}(t))$
[…]
$:=\bm{I}_{1}(t)+\bm{I}_{2}(t),$
where the second equality is from fundamental theorem of calculus and $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)=\bm{w}(t)-s\eta\bm{J}(t)^{T}\bm{H}(t)^{-1}\bm{u}(t)$.
Note that $\frac{\partial\bm{u}(\bm{w}(t))}{\partial\bm{w}}=\bm{J}(t)$, thus $\bm{I}_{1}(t)=\eta\bm{u}(t)$. Plugging it into (85) yields that
$\bm{u}(t+1)=(1-\eta)\bm{u}(t)+\bm{I}_{2}(t).$ (87)
Therefore, it remains only to bound $\|\bm{I}_{2}(t)\|_{2}$.
[…]
$\lesssim\frac{\eta\sqrt{L(t)}}{\sqrt{\lambda_{0}}}R^{{}^{\prime\prime}}$
where the last inequality follows from the fact that
$\|\bm{w}_{r}(s)-\bm{w}_{r}(0)\|_{2}\leq s\|\bm{w}_{r}(t+1)-\bm{w}_{r}(0)\|_{2}+(1-s)\|\bm{w}_{r}(t)-\bm{w}_{r}(0)\|_{2}\leq R^{{}^{\prime}}\leq R$
and Lemma 7.
Plugging (88) into the recursion formula (87) yields that
$\|\bm{u}(t+1)\|_{2}^{2}$ $=\|(1-\eta)\bm{u}(t)+\bm{I}_{2}(t)\|_{2}^{2}$
$=(1-\eta)^{2}\|\bm{u}(t)\|_{2}^{2}+\|\bm{I}_{2}(t)\|_{2}^{2}+2\langle(1-\eta)\bm{u}(t),\bm{I}_{2}(t)\rangle$
$\leq(1-\eta)^{2}\|\bm{u}(t)\|_{2}^{2}+\|\bm{I}_{2}(t)\|_{2}^{2}+2(1-\eta)\|\bm{u}(t)\|_{2}\|\bm{I}_{2}(t)\|_{2}$
$\leq\left[(1-\eta)^{2}+\frac{C^{2}\eta^{2}(R^{{}^{\prime\prime}})^{2}}{\lambda_{0}}+2(1-\eta)\frac{C\eta R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\right]\|u(t)\|_{2}^{2}$
$\leq(1-\eta)\|\bm{u}(t)\|_{2}^{2},$
where the last inequality follows from the requirement
$2\frac{C\eta R^{{}^{\prime\prime}}}{\sqrt{\lambda_{0}}}\leq\frac{\eta}{4},$
i.e., $R^{{}^{\prime\prime}}\lesssim\sqrt{\lambda_{0}}$.
Therefore, we can deduce that $\|\bm{u}(t)\|_{2}^{2}\leq(1-\eta)^{t}\|\bm{u}(0)\|_{2}^{2}$ holds for $t=0,\cdots,k$.
8.3. Proof of Corollary 1
Proof.
Setting $\eta=1$ in (85) yields that
$\bm{u}(t+1)=\bm{I}_{2}(t).$
From (88), we have that
[…]
As $\bm{w}(s)=s\bm{w}(t+1)+(1-s)\bm{w}(t)$, then for any $r\in[m]$, we have $\|\bm{w}_{r}(s)\|_{2}\leq s\|\bm{w}_{r}(t+1)\|_{2}+(1-s)\|\bm{w}_{r}(t)\|_{2}\leq B$.
When $\sigma$ is smooth, we can deduce that for any $r\in[m]$,
$\left\|\frac{\partial s_{p}(\bm{w}(s))}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(\bm{w}(t))}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{1}m}}(B^{2}+1)\|\bm{w}_{r}(s)-\bm{w}_{r}(t)\|_{2}\leq\frac{1}{\sqrt{n_{1}m}}(B^{2}+1)\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}$
and
$\left\|\frac{\partial h_{j}(\bm{w}(s))}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(\bm{w}(t))}{\partial\bm{w}_{r}}\right\|_{2}\lesssim\frac{1}{\sqrt{n_{1}m}}(B+1)\|\bm{w}_{r}(s)-\bm{w}_{r}(t)\|_{2}\leq\frac{1}{\sqrt{n_{1}m}}(B+1)\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}.$
From (85), we know that for any $r\in[m]$,
$\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}\lesssim\frac{B^{2}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}.$
Thus for any $s\in[0,1]$, we have
$\|\bm{J}(\bm{w}(s))-\bm{J}(\bm{w}(t))\|_{2}^{2}$
$\leq\sum\limits_{r=1}^{m}\left(\sum\limits_{p=1}^{n_{1}}\left\|\frac{\partial s_{p}(\bm{w}(s))}{\partial\bm{w}_{r}}-\frac{\partial s_{p}(\bm{w}(t))}{\partial\bm{w}_{r}}\right\|_{2}^{2}+\left\|\frac{\partial h_{j}(\bm{w}(s))}{\partial\bm{w}_{r}}-\frac{\partial h_{j}(\bm{w}(t))}{\partial\bm{w}_{r}}\right\|_{2}^{2}\right)$
$\lesssim\frac{1}{m}\sum\limits_{r=1}^{m}\left((B^{4}+1)\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}^{2}+(B^{2}+1)\|\bm{w}_{r}(t+1)-\bm{w}_{r}(t)\|_{2}^{2}\right)$
$\lesssim B^{4}\left(\frac{B^{2}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}\right)^{2}.$
Plugging it into (89), we have
[…]
$\lesssim\frac{\sqrt{L(t)}}{\sqrt{\lambda_{0}}}\frac{\eta B^{4}}{\sqrt{m}\lambda_{0}}\sqrt{L(t)}$
$=\frac{B^{4}}{\sqrt{m\lambda_{0}^{3}}}L(t).$
Combining with the fact $\bm{u}(t+1)=\bm{I}_{2}(t)$ yields that
$\left\|\begin{pmatrix}\bm{s}(t+1)\\ \bm{h}(t+1)\end{pmatrix}\right\|_{2}\leq\frac{CB^{4}}{\sqrt{m\lambda_{0}^{3}}}\left\|\begin{pmatrix}\bm{s}(t)\\ \bm{h}(t)\end{pmatrix}\right\|_{2}^{2}$
holds for $t\in\mathbb{N}$, where $C$ is a universal constant. ∎
9. Auxiliary Lemmas
Lemma 8 (Theorem 3.1 in [24]).
If $X_{1},\cdots,X_{n}$ are independent mean zero random variables with $\|X_{i}\|_{\psi_{\alpha}}<\infty$ for all $1\leq i\leq n$ and some $\alpha>0$, then for any vector $a=(a_{1},\cdots,a_{n})\in\mathbb{R}^{n}$, the following holds ture:
$P(|\sum\limits_{i=1}^{n}a_{i}X_{i}|\geq 2eC(\alpha)\|b\|_{2}\sqrt{t}+2eL_{n}^{*}(\alpha)t^{1/\alpha}\|b\|_{\beta(\alpha)})\leq 2e^{-t},\ for\ all\ t\geq 0,$
where $b=(a_{1}\|X_{1}\|_{\psi_{\alpha}},\cdots,a_{n}\|X_{n}\|_{\psi_{\alpha}})\in\mathbb{R}^{n}$,
$C(\alpha):=\max\{\sqrt{2},2^{1/\alpha}\}\left\{\begin{aligned} \sqrt{8}(2\pi)^{1/4}e^{1/24}(e^{2/e}/\alpha)^{1/\alpha}&,&if\ \alpha<1,\\ 4e+2(\log 2)^{1/\alpha}&,&if\ \alpha\geq 1.\end{aligned}\right.$
and for $\beta(\alpha)=\infty$ when $\alpha\geq 1$ and $\beta(\alpha)=\alpha/(\alpha-1)$ when $\alpha>1$,
$L_{n}(\alpha):=\frac{4^{1/\alpha}}{\sqrt{2}\|b\|_{2}}\times\left\{\begin{aligned} &\|b\|_{\beta(\alpha)},&if\ \alpha<1,\\ &4e\|b\|_{\beta(\alpha)}/C(\alpha),&if\alpha\geq 1.\end{aligned}\right.$
and $L^{*}_{n}(\alpha)=L_{n}(\alpha)C(\alpha)\|b\|_{2}/\|b\|_{\beta(\alpha)}$.
Lemma 9 (Bernstein inequality, Theorem 3.1.7 in [25]).
Let $X_{i}$, $1\leq i\leq n$ be independent centered random variables a.s. bounded by $c<\infty$ in absolute value. Set $\sigma^{2}=1/n\sum_{i=1}^{n}\mathbb{E}X_{i}^{2}$ and $S_{n}=1/n\sum_{i=1}^{n}X_{i}$. Then, for all $t\geq 0$,
$P\left(S_{n}\geq\sqrt{\frac{2\sigma^{2}t}{n}}+\frac{ct}{3n}\right)\leq e^{-u}.$
Lemma 10.
For $10<\delta<$ 1, with probability at least $1-\delta$, we have
$\|\bm{y}-\bm{u}(0)\|_{2}^{2}=\mathcal{O}\left(n\log\left(\frac{n}{\delta}\right)\right).$
Proof.
First, from Cauchy’s inequality, we have
$\|\bm{y}-\bm{u}(0)\|_{2}^{2}$ $=\sum\limits_{i=1}^{n}\left(y_{i}-\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}$
$\leq\sum\limits_{i=1}^{n}2y_{i}^{2}+2\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}$
$=2\sum\limits_{i=1}^{n}y_{i}^{2}+2\sum\limits_{i=1}^{n}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right)^{2}.$
Since $\bm{w}_{r}(0)^{T}\bm{x}_{i}\sim\mathcal{N}(0,\|\bm{x}_{i}\|_{2})$ and $\|\bm{x}_{i}\|_{2}\leq\sqrt{2}$, we can deduce that
$\|a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\|_{\psi_{2}}\leq\|\bm{w}_{r}(0)^{T}\bm{x}_{i}\|_{\psi_{2}}=\mathcal{O}(1)$
holds for all $r\in[m]$ and $i\in[n]$.
Thus for any fixed $i\in[n]$, applying Lemma 8 yields that with probability at least $1-2e^{-t}$,
$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{x}_{i})\right|\leq C\sqrt{t},$
where $C$ is a universal constant. By taking union bound on all $i\in[n]$, the above inequality holds for all $i\in[n]$ with probability at least $1-2ne^{-t}$.
Let $2ne^{-t}=\delta$, i.e., $t=\log\left(\frac{2n}{\delta}\right)$, then we have
$\|\bm{y}-\bm{u}(0)\|_{2}^{2}\lesssim n+nt\lesssim n\log\left(\frac{n}{\delta}\right).$
Lemma 11.
For $10<\delta<$ 1, with probability at least $1-\delta$, we have that when $m\geq\log^{2}\left(\frac{n_{1}+n_{2}}{\delta}\right)$,
$L(0)=\left\|\begin{pmatrix}\bm{s}(0)\\ \bm{h}(0)\end{pmatrix}\right\|_{2}^{2}=\mathcal{O}\left(d^{6}\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$
Proof.
Recall that for $p\in[n_{1}]$,
$s_{p}(0)=\frac{1}{\sqrt{n_{1}}}\left[\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)-f(x_{p})\right]$
and for $j\in[n_{2}]$,
$h_{j}(0)=\frac{1}{\sqrt{n_{2}}}\left[\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})-g(\bm{y}_{j})\right].$
Then
$L(0)$ $=\sum\limits_{p=1}^{n_{1}}\frac{1}{2}(s_{p}(0))^{2}+\sum\limits_{j=1}^{n_{2}}\frac{1}{2}(h_{j}(0))^{2}$
$\leq\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right)+\frac{1}{n_{1}}\sum\limits_{p=1}^{n_{1}}f^{2}(x_{p})$
$\quad+\frac{1}{n_{2}}\sum\limits_{j=1}^{n_{2}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})\right)^{2}+\frac{1}{n_{2}}\sum\limits_{j=1}^{n_{2}}g^{2}(\bm{y}_{j}).$
Note that
$|a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)|\lesssim\|\bm{w}_{r}(0)\|_{2}^{3}$
and $|a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})|\lesssim\|\bm{w}_{r}(0)\|_{2}^{3}.$
Since $\|\|\bm{w}_{r}(0)\|_{2}^{3}\|_{\psi_{\frac{2}{3}}}\leq(\|\|\bm{w}_{r}(0)\|_{2}^{2}\|_{\psi_{1}})^{\frac{3}{2}}\lesssim d^{3}$, from Lemma 8, we have that for fixed $i\in[n_{1}]$ and $j\in[n_{2}]$ with probability at least $1-2e^{-t}$,
$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\left(\sigma^{{}^{\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})w_{r0}(0)-\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(0)^{T}\bm{x}_{p})\|\bm{w}_{r1}(0)\|_{2}^{2}\right)\right|\lesssim d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}}$
and with probability at least $1-2e^{-t}$,
$\left|\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma(\bm{w}_{r}(0)^{T}\bm{y}_{j})\right|\lesssim d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}}.$
Then taking a union bound for all $i\in[n_{1}]$ and $j\in[n_{2}]$ with $2(n_{1}+n_{2})e^{-t}=\delta$ yields that
$L(0)$ $\lesssim(d^{3}\sqrt{t}+\frac{d^{3}}{\sqrt{m}}t^{\frac{3}{2}})^{2}$
$\lesssim d^{6}t+\frac{d^{6}t^{3}}{m}$
$=d^{6}\left(\log\left(\frac{n_{1}+n_{2}}{\delta}\right)+\frac{1}{m}\log^{3}\left(\frac{n_{1}+n_{2}}{\delta}\right)\right)$
$\lesssim d^{6}\log\left(\frac{n_{1}+n_{2}}{\delta}\right),$
as $m\geq\log^{2}\left(\frac{n_{1}+n_{2}}{\delta}\right)$.